### v2

# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

Abstract
The growing prevalence of generative AI inevitably raises concerns regarding the associated risks and safety implications, which catalyzes significant progress in AI safety. However, as this field thrives, a critical question emerges: Are our current efforts aligned with the broader perspective of human history and civilization? This paper presents a blueprint for an advanced human society and leverages this vision to guide contemporary AI safety efforts. It outlines a future where the Internet of Everything becomes reality, and create a roadmap of significant technological advancements towards this envisioned future. For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face. By projecting current efforts against this blueprint, we examine the alignment between the present efforts and the long-term needs. This paper identifies gaps in current approaches and highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible future for AI development. This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby fostering AI’s role in promoting a more secure and sustainable future for human civilization.
1 Introduction
The rapid development of AI and Large Language Models (LLMs) has fostered extensive progress in AI safety, specifically, LLM safety [171, 143, 51, 230, 90, 37]. Researchers have been dedicated to addressing potential security and privacy risks in LLM lifecycle, aiming to align AI behaviors with human values and prevent misuse of LLMs, inappropriate outputs, information leakage, etc. However, despite the intense focus on LLM safety, a critical question emerges: Are current AI safety efforts truly aligned with the long-term evolution of human civilization, or are they simply addressing the concerns of today?
One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [146]. Despite their impressive capabilities in natural language processing and problem-solving [270, 265, 54, 38], today’s AI, including advanced LLMs [286, 260, 231], falls short of what could be considered genuine intelligence. These systems heavily rely on vast training datasets to function effectively, and lack true understanding, consciousness, self-awareness, and the ability to reason in ways comparable to human cognition. They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities capable of abstract thought and genuine reasoning. Moreover, recent studies have raised concerns about the reasoning abilities of LLMs, which may actually be a form of approximate retrieval from their training data and the deductive closure of the training data [96, 239]. While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it fundamentally differs from human reasoning that involves abstract thinking, causal understanding, and the ability to generalize from limited examples.
The unresolved issue of sustainable energy is another critical factor that challenges our current efforts to AI safety [5]. Current AI fails to fully represent a third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved. Historically, the two industrial revolutions that have profoundly shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by a new wave of technological innovations powered by electricity and petroleum [76]. However, current AI, rather than solving energy issues, consumes a significant amount of energy. Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent and approximately 0.02% of California’s annual electricity generation [7, 39, 169]. Also, inference with LLMs is computationally intense, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [140, 72]. Given that ChatGPT alone has over 200 million weekly active users and receives over 1.54 billion page visits monthly [213], when scaled to billions of queries, the energy consumption becomes substantial. The significant energy requirements of current AI systems raise serious sustainability concerns and limiting their deployment throughout the human society. Until we overcome the energy constraints, AI’s ability to reshape society remains restricted.
Figure 1: Illustration of Internet of Everything in a Household
These concerns suggest that our efforts on AI safety today may fail to capture deeper insights that could emerge as AI evolves, potentially neglecting technological challenges that ultimately shape the impact of AI on human civilization. This raises several critical questions:
  • Are we focusing on the right problems that align with the long-term AI development?
  • Are we potentially overlooking significant challenges that will emerge as AI evolves?
  • Are we on the right track for addressing real challenges in AI safety?
“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety research. In this advanced future, energy issues have been resolved, and intelligent chips and brain-computer interfaces have been fully developed, which enables advanced intelligent robots to integrate into everyday life. Breakthroughs in foundational theories, such as Einstein’s relativity [57, 58] and Quantum Mechanism [59, 279], have catalyzed revolutionary progress in AI, and humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1. While this vision may be beyond the horizon of several decades, by reviewing the present landscape of AI safety against this far future, we can gain deeper insights into the potential and limitations of our current approaches, as well as specific missions for AI safety researchers in the 2020s. It reminds us that, while addressing today’s immediate concerns, we should avoid being constrained by these short-term solutions and look toward the future to guide our present efforts.
Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” to encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape. ii) Deeper collaboration with experts from diverse fields is essential to align AI with human values better. iii) AI workflow customization and user-AI interactions are important, especially in dynamic contexts and across different user groups.
Humanity’s pursuit of higher technology will never stop. While searching for the sixpence on the ground, do not forget to look up at the night sky and look for the moon.
Paper Type Topic Notes
“Grounding and Evaluation for LLMs” [102] Survey Grounding; safety Offer valuable insights for enterprises
“Current state of risks” [8] Survey Guardrails Aspects of guardrails and challenges of implementations
“Exploring Vulnerabilities and Protections” [122] Survey Vulnerabilities Vulnerabilities of LLMs, especially prompt injections and backdoor; defenses against the attacks
“AI safety in LLMs” [37] Survey Safety issues in LLM lifecycle Safety issues towards entities in LLM lifecycle; Data/model/prompt safety; alignment; safety at scale
“Security risk” [185] Survey Attacks Attacks and risks in AI models; defensive methods
ByteDance alignment [124] Survey Broad concepts Reliability, fairness, explainability, etc
“Security of AI Agents” [82] Survey Agent security Security issues, vulnerabilities, and defenses for agents
“The good, the bad, the ugly” [272] Survey Risks; vulnerabilities; defenses LLMs as tools for security and privacy; potential risks and threats; vulnerabilities and defenses.
“Use of LLMs” [142] Survey Vulnerabilities, threats, defenses, Prevention measures, vulnerabilities, and methods to mitigate the risks associated with the misuse of LLMs
“AI risk management” [181] Insight Safety; security; governance Bridging AI safety and security in risk management; introducing a reference framework to facilitate common understanding of their differences and interplay.
“Human-ai safety” [11] Insight The integration of AI and control systems safety A unifying framework to capture safety-critical human–AI interactions; a roadmap towards next-generation human-centered AI safety
“Managing AI risks” [13] Insight Governance Emphasizing the urgent need for a balanced approach combining technical research and proactive governance to manage the extreme risks of AI systems
“Guaranteed safe AI” [43] Position Framework Define Guaranteed Safe AI to equip AI systems with formal, verifiable and auditable safety guarantees.
DeepMind ethics paper [67] Vision AI ethics Ethical challenges; interactions; impacts on society
DeepMind evaluation [252] Insight Safety evaluation Safety evaluation for genAI systems at DeepMind
Safety practices at OpenAI [151, 152, 154, 155] Blog Safety practices; framework “Lessons learned” [151], “Reimagining secure infrastructure” [152], “Preparedness Framework” [154], “OpenAI safety update” [155]
“AI governance comprehensive” [214] Book Governance AI governance framework; Case studies in industries
Table 1: Overview of AI Safety Papers
2 Related Works
Multiple AI safety surveys and benchmarks have been published in recent years. These studies focus on different aspects of AI safety, such as jailbreaking, red teaming, guardrail, ethics, etc. A summary of these works is presented in Table 1, which provides an overview of the related works and their focus within the field of AI safety.
Among these works, [124] provides comprehensive insights into building trustworthy AI systems by focusing on reliability, safety, fairness, robustness, and adherence to social norms. [37] addresses AI safety concerns from the perspective of different entities, e.g., data and prompts, in LLM lifecycle, and discusses data safety, model safety, prompt safety, alignment, and the complexities of scaling.  [102] provides valuable enterprise insights into the practical challenges and opportunities for AI safety in industries. [67] provides a systematic analysis of the ethical risks of advanced AI assistants, and discusses their potential influence on users, the risks of misuse, strategies for enhancing human-AI interactions, and the broader implications for human society. [43] proposes the concept of “Guaranteed Safe” (GS) AI, emphasizing the need for formal, verifiable, and auditable safety guarantees to ensure robust AI behavior.
[11] proposes a technical roadmap towards next-generation human-centered AI safety that aims to anticipate and prevent potential hazards in the interactions between AI systems and users. Specifically, the paper bridges the fields of AI safety and control systems, and highlights the open challenges at their intersection. The authors emphasize the importance of leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for achieving meaningful safety assurances.
Different from existing works, this paper envisions an advanced society in the future of human civilization, and use this vision to guide current AI safety research. By comparing the current landscape of AI safety with this distant future, this work identifies both the potential and limitations of today’s approaches and suggests specific missions for AI safety researchers in the 2020s. This perspective encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
3 From the New World
In this future, humanity has entered an unprecedented era. Scientists have achieved breakthroughs in energy production with technological innovations. The resulting flood of abundant, clean energy has revolutionized human civilization beyond the our ancestors’ wildest dreams, reshaping the world and initiating a new chapter of human civilization.
3.1 Imagine the Future: A Blueprint
The transformative energy revolution has catalyzed human civilization from every aspect. Foundational theories of science have undergone revolutionary breakthroughs, with major advancements in nanotechnology [136, 19, 61], Quantum Mechanics [59, 279], and Einstein’s relativity theories [57, 58]. The world has been reshaped by innovative technologies such as intelligent chips [86, 115], brain-computer interfaces [147, 81, 254], holographic technology [22, 273, 15], and advanced 3D printing [205, 73], bringing changes in daily life of miraculous.
Transportation has evolved dramatically. While simple mechanical vehicles, such as fixed-track trains and high-speed rail [144, 179, 109], still exist due to their efficient and straightforward design, they operate at astonishing speeds with high stability, powered by efficient and limitless energy. Advanced autonomous vehicles and aircraft are widespread and are controlled by intelligent chips, eliminating the need for any human intervention. Also, due to breakthroughs in Einstein’s relativity theories [57, 58] and Quantum Mechanics [59, 279], humanity has unlocked the potential of instant transportation that allow for long-distance travel in short time that transcends the imagination of the 21st century.
Advanced brain-computer interfaces [147, 81, 254] and intelligent microchips [86, 115] has revolutionized the way people perceive, which expands the boundaries of human cognition. No longer are people operating devices and asking questions in search engines or LLMs; the embedded intelligence chips help them to gather more comprehensive information and interact with the surrounding environment. Meanwhile, communication over long distances has transcended the limitations of cables and the outdated internet; information, thoughts, emotions, and sensations flow through a new medium due to the advancement in quantum computing [77, 198].
Intelligent robots have been integrated into every aspect of modern society and play important roles of daily life; see Figure 1. Different from their “data-hungry machine intelligent ancestors” that relies on huge amount of training data and statistical patterns in the early 21st century [131, 286], these intelligent robots perceive and learn through intelligent chips, which are real artificial intelligence and possess genuine understanding and adaptability, much like a human child discovering the world for the first time. Also, their functionalities are customizable through specialized chips based on user needs. Robots can engage in complex interactions with people and adapt their capabilities and expertise to suit their human counterparts’ requirements.
In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds have been blurred, i.e., Internet of Everything. Such a network integrates objects with embedded intelligent chips, such as smart devices, autonomous systems, intelligent robots, and humans cognition, which allows for unprecedented levels of efficiency, personalization, and adaptability and fundamentally reshapes the way people interact with the world.
3.2 Retrospective on AI Safety in the New World
In a world where the Internet of Everything has become a reality, intelligent robots are customized to meet diverse user needs and seamlessly integrate into daily life; see Figure 1. Consequently, the concept of AI safety has been redefined to focus more on the quality of services they provide, i.e., emphasizing comprehensive quality assurance. Unlike today’s command-and-response interactions between humans and LLMs, future robots can engage in sophisticated reasoning, emotional cognition, and possess appropriate levels of autonomy. They will also interact dynamically with other entities, including humans and fellow robots, within the interconnected system.
The Internet of Everything scenarios necessitate the robots to handle complex interactions, demanding a sophisticated understanding of the environment for decision-making. Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level. Such advanced interactions require deep understanding of the environments, demanding real-time information retrieval and processing, which is far beyond what can be achieved with simple code logic. Robots must continuously collect and analyze data from their environment, and integrate external inputs with contextual knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align precisely with human needs and expectations.
Customization is a key aspect of future robots for providing personalized services to meet individuals’ requirements. Even robots of the same type should be capable of adapting their services to different individuals. Teaching robots should adjust educational materials and teaching methods to capitalize on each student’s unique talents and strengths, while cleaning robots should intelligently schedule their works to minimize disruption to other humans in the household, e.g., avoiding occupied workspace. Such customization aligns the robots’ capabilities with specific user preferences, and ensures that the robots serve as effective assistants across a wide range of daily tasks and environments.
Finally, the effectiveness of intelligent robots requires a balance between robot autonomy with human oversights. Too little autonomy would burden users with continuous supervision, while too much autonomy could raise safety concerns or result in unintended behaviors. For instance, a care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions. Conversely, while the tasks of cleaning robots are generally routine, they must still adapt to specific circumstances, such as recognizing when an area is occupied by people and postponing cleaning activities. An appropriate level of robot autonomy optimizes the division of labor between humans and robots, ensuring that robots function effectively and safely while enhancing our daily lives without compromising human control over these intelligent robots.
4 Bridging Today and the Future
Figure 2: Evolution of Human Civilization
This section presents an ambitious roadmap towards the aspirations for the advanced future, exploring key milestones in the evolution of AI and human civilization. By considering potential paradigm shifts in fundamental theories and breakthroughs in technology, we feature this fascinating journey into three eras, i.e., i) the data-driven era where AI learns knowledge from vast training datasets, ii) the neural augmentation era where human-machine interface [147, 81, 254] and intelligent chips [86, 115] extend the boundary of human cognition, and iii) the revolutionary era with a breakthrough in energy production that catalyzes transformative changes across every aspect of human society; see §4.1, §4.2, and §4.3, respectively. We then examine the evolving paradigms of AI safety throughout this journey, analyzing the potential focuses and challenges of each era and exploring the impacts of technological advancement on AI safety in §$4.$
4.1 Rookie Level: Data-Driven Era
The Data-Driven Era represents the early phase of AI development, where AI models heavily rely on the quality, availability, and quantity of training data [131]. AI makes probabilistic inferences based on patterns found in massive training datasets, thus failing to achieve genuine understanding of data [96, 239]. While AI in this era cannot be regarded as real intelligence, it assists people to solve problems and enhances productivity in the society.
Stage 1: Past, Current, and the Near Future. This stage is characterized by diverse AI models for processing various forms of data across different scenarios [131, 50, 63, 210, 231, 260]. Researchers have been engaging in refining model structures and developing advanced algorithms, enabling more sophisticated analysis of data in different formats. The thrive of LLMs exemplifies this progress, showcasing unprecedented capabilities of AI models in language processing and generation [260, 286, 231]. Concurrently, we observe significant advancements in other fields such as computer vision [101, 221, 256], with progress in image processing [235, 25, 216] and video analysis [223, 261, 267], as well as advancements in audio recognition [64, 178, 4, 106]. This stage is driven by exploring vast datasets and the refining algorithms continuously, which consistently pushes the boundaries of AI capabilities within specific domains.
Stage 2: Interdisciplinary AI. This stage is characterized by leveraging AI to transform traditional fields, such as healthcare [231, 2], finance [260, 282], and education [184, 135], through cross-disciplinary applications in human society. Mature multi-agent systems [116, 246, 263, 79] that leverage AI entities with different specializations will be widely utilized to assist people and solve sophisticated problems. Such agents participate in decision-making processes, or represent users to finish some tasks, providing more sophisticated solutions of a given problem. In healthcare, for example, a multi-agent system may involve collaboration between various medical agent specialties and integrate data from diverse sources to formulate comprehensive treatment plans. The rise of interdisciplinary AI integration will amplify the utility of AI in more complex scenarios. Although AI at this stage is still not real intelligence, it becomes deeply integrated in different domains in human society, and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
4.2 Champion Level: Neural Augmentation Era
In this era, AI models have been widely applied in human society, and humans may witness breakthroughs in advanced technologies and fundamental theories, which may extend the boundary of human cognition and life experience.
Stage 3: Breakthroughs in Advanced Technologies. In this stage, breakthroughs in advanced technologies emerge, such as brain-computer interfaces [147, 81, 254], nanotechnology [136, 19, 61], and holographic technology [22, 273, 15]. These innovations enable unprecedented levels of interaction between humans and machines by augmenting human senses and cognition. For instance, intelligent chips [86, 115] may be developed that integrate directly with the human body, creating a seamless interface between biological and digital systems. Such advances also change the nature of the data. Unlike traditional static formats such as text, images, or videos, the data in this stage becomes more comprehensive, encompassing dynamic, real-time, and sensory-rich formats, such as neural activity patterns or real-time environmental data from augmented reality systems [31, 32]. In addition to the advancements in AI seen during the Data-Driven Era, this stage further expands the boundaries of human perception and cognition, revolutionizing how people experience and interact with the world. However, widespread application of these techniques is limited by energy constraints and the significant resources required for implementation.
Stage 4: Breakthroughs in Fundamental Theories. In this stage, humanity experiences a new wave of breakthroughs in fundamental theories, such as Einstein’s Relativity Theories [57, 58] and Quantum Mechanics [59, 279]. These theoretical advances redefine our understanding of the universe and offer new possibilities for innovations in technologies. Such breakthroughs may also unlock solutions to longstanding challenges in human history, i.e., the energy issues, offering clean, abundant, and sustainable energy sources.
4.3 Ultimate Level: Revolutionary Era
Stage 5: Energy Revolution. A new era arrives as humanity finally achieves energy revolution, unraveling the mysteries of sustainable power generation. This may be achieved through mature technologies that are currently on the horizon, such as breakthroughs in large-scale controlled nuclear fusion [70, 200, 183], or revolutionary solar technologies [23, 137], together with advancement in nanotechnologies [136, 19, 61] that produce new materials for collecting and storing solar energy efficiently. The energy issue may also be solved through innovations that are beyond the scope of today’s technologies, such as deploying satellites for gathering and storing solar energy in space. These satellites might then transmit the abundant energy to the receivers on Earth. Regardless of the specific technology, the flood of clean, abundant energy can eliminate the dependence on fossil fuels, and fundamentally revolutionize the global economic landscapes as well as human civilization.
Stage 6: The Internet of Everything. The final stage is the realization of Internet of Everything. Every object, living being, and system becomes part of a vast, intelligent network. This hyper-connected world is characterized by unprecedented levels of perception, communication, and interactions between humans and intelligent robots in an area. Language barriers may disappear, as human can perceive with chips for understanding, which enables ubiquitous real-time context-aware translations between people of different languages and cultures. Due to everyone is connected to a huge network, the boundary of human sense has been extended, which may also catalyze breakthroughs in holographic technology [15, 22, 273]. Remote experiences become true; working, entertainment, and communication are revolutionized, and human has entered the ultimate advanced era.
Figure 3: Production forces of AI & core safety challenges in different historical stages.
4.4 Evolving AI Safety Paradigms
The landscape of AI safety shifts as humanity processes through stages of technology revolutions, and the focus of AI safety is different at each stage. The core AI safety challenges in different historical eras are summarized in Figure 3.
Data-Driven Era. Current AI safety practices primarily focus on data safety and model safety [37, 102, 42, 220, 167, 193, 253, 225, 224]. A critical aspect involves data privacy and security, particularly given the vast amounts of sensitive information used in AI training [118, 120, 83, 117]. AI safety researchers and practitioners also explore adversarial attacks to training data or AI models [9, 35, 119, 244, 211, 269], as well as malicious inputs designed to manipulate AI systems or induce producing unsafe content, e.g., jailbreaking or prompt injection attacks [75, 292, 126, 33, 90, 258, 293]. Correspondingly, red-teaming ML models [12, 68, 173, 197, 243] is widely used to identify potential vulnerabilities in ML models with attacks to further enhance the robustness of AI models, and significant work has been proposed to enhance security and robustness of AI systems [27, 240, 290]. Additionally, researchers have been engaging in AI governance to enhance AI accountability [21], transparency and explainability [291] in AI systems to ensure the decision-making processes are understandable, accountable, and aligned with regulations. Other significant areas of focus is aligning AI outputs with human values [249, 174, 232, 206, 245].
Neural Augmentation Era. Advancements in cognitive chips and brain-computer interfaces have extended the boundaries of human cognition and have changed the way of human perceiving the world. With the integration of intelligent chips and their direct neural connections to AI systems, users can interact with AI with high flexibility, and the customization of AI services towards different user groups or domain knowledge might be a critical focus [248]. Individuals might rely on AI in different ways based on their background and cognitive needs, thus, personalized experiences of AI services should be explored for different user needs and preferences.
Revolutionary Era. In the revolutionary era, AI safety challenges may focus on the complex interactions between interconnected entities, such as humans and intelligent robots, within the network. As users may have diverse needs and robots will provide customized services, ensuring the safety and reliability of these interactions becomes challenging. Robots may need to request and integrate data from multiple external sources in real-time to fulfill user requests. Managing the quality and security of these interactions is crucial, as any breakdown or interference could lead to incorrect or even harmful outcomes. Ensuring seamless communication, data integrity, and decision-making in such a sophisticated network may require smarter, more efficient methods, such as advanced cache management systems, to support the dynamic and complex interactions. This requires addressing issues of latency, data accuracy, and the security of transmitted information to maintain a safe and reliable network to ensure high-quality interactions between humans and robots in the network.
By aligning AI with the progress of human civilization, we can see the path of AI safety evolves in different stages, which reflects the growing complexity of the roles of AI in human society. Initially, the focus is on addressing safety issues related to data and models. As AI and novel technologies become more developed, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs. Finally, when entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the same network.
These insights offer guidelines for what we can do for AI safety at the current stage. In the following sections, we will overview today’s AI safety landscape, then reflect on the progress in this area so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed in the present stage.
5 Back to the 2020s: Today’s AI Safety Landscape
This section overviews mainstream AI safety practices so far, and project them against the blueprint future in §3.1 to guide our current efforts in AI safety.
5.1 Jailbreaking and Red Teaming
Jailbreaking [49, 123, 175, 207, 215, 219, 234, 250, 258, 268, 293] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs. Malicious users or adversaries utilize these techniques to manipulate the model into producing abnormal or harmful responses, such as giving instructions for illegal activities and generating offensive content. Common jailbreak methods include crafting inputs to trick the model, planting backdoors to exploiting vulnerabilities of the model, etc. Such activities pose significant risks by undermining safety measures of AI models, making it challenging to ensure the model to behaves appropriately across all scenarios.
Red-teaming is a common practice for mitigating unsafe behaviors in LLMs, where teams of experts simulate attacks or misuse scenarios to identify potential vulnerabilities in AI models [68, 243, 197, 12, 173, 128, 275]. By actively challenging AI models and probing for weaknesses through adversarial inputs, bias exploitation, or system manipulation, it helps uncover potential risks that may not be evident in normal usage, such as biases, or failure modes that might lead to undesirable behaviors in AI models. The outputs from red-teaming can be used to train the model to be less likely to cause harm or steer it away from undesirable outputs. This proactive approach has proven effective in stress-testing AI models and has been widely used in industry practices [18, 173, 68].
As AI systems grow more sophisticated, red teaming must continually adapt to keep pace with increasingly complex attacks, which calls for more advanced tools and interdisciplinary expertise [153]. Also, model owners can conduct regular and diverse red teaming to anticipate and mitigate risks in evolving use cases. Moreover, current red-teaming practices are often static, typically conducted after model pretraining and consisting of one-time attacks. While this approach provides useful insights, it does not fully account for the evolving user needs, especially when dealing with customized AI services. To address this, red teaming should simulate diverse user groups and scenarios, such as simulating a diverse range of user groups and their unique needs, to address the customization of AI services and enable a comprehensive evaluation of potential vulnerabilities across different contexts.
5.2 Content Moderation and Guardrails
Content moderation involves detecting and reviewing user input and AI-generated content to ensure it complies with safety, ethical, and regulatory standards [133, 112, 80, 91]. Moderation approaches leverages rule-based methods, machine learning classifiers, and human oversight to monitor, evaluate, and manage inputs and outputs of AI models. These approaches identify and address inappropriate content, such as toxicity, bias, hallucinations, sensitive information, and jailbreaks, to ensure that the AI models do not produce or engage with harmful content, which maintains responsible AI behaviors.
Guardrails [141, 189] employ customizable AI workflows for ensuring the quality of the model outputs while enhancing the control of the whole workflows. It functions as an intermediary layer between users and AI models, enabling users to add components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
As we reflect on practices for AI moderation and guardrails, several key challenges and considerations have emerged. The balance between utility and safety is crucial and may require dynamic solutions according to the changing contexts. Overly strict moderation can cause an AI system to be helpless, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless. Thus, defining appropriate thresholds for acceptable responses is challenging and requires careful consideration of the changing context and potential risks. Also, customization for moderation should be applied for different use cases and user groups, as certain information may be safe for one user group but inappropriate or dangerous for another. As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public.
Topic Research Industry Practices
Jailbreaking [90, 33, 292, 258, 293, 126, 125, 170, 36, 177, 175, 219, 234, 123, 250, 49, 207, 215, 268], etc PromptArmor [227]
Red teaming [68, 243, 197, 12, 173, 128, 275, 230, 134, 48], etc Purple Llama [18], DeepMind [173], Anthropic [68], Protect AI [220], Giskard [42], Virtue AI [228], Dynamo AI [55], Mindgard [294]
Moderation [108, 259, 182, 60, 130, 92, 172, 266, 34, 107], etc OpenAI Moderation API [133], Perspective API [112], Detoxify [80], Llama Guard [91], Emergence AI [148], Giskard [42], Dynamo AI [55], Calypso AI [26], Lakera AI [226], BreezeML [24]
Guardrail - Guardrails AI [141], Nvidia Nemo Guardrails [189], Protect AI [220]
Privacy [53, 118, 139, 30, 288, 87, 281, 204], etc Dynamo AI [55], ProtectAI [220], Private AI [69], PromptArmor [227], DataGrail [193], Dastra [225], OneTrust [224], Relyance AI [99], Zendata [229]
Defense Methods [199, 88, 240, 176, 257, 262, 133, 29, 66, 247, 278, 88, 168, 74, 40, 51, 209, 208], etc OpenAI Rule-based rewards [143]; DeepMind Weight Averaged Reward Models [188]; others see Moderation, Guardrail, and Privacy
Serving Security [258] ProtectAI [220], Relyance AI [99], Transcend [100]
AI governance [276, 277, 6, 104, 56], etc Giskard [42], ProtectAI [220], Calypso AI [26], Saidot [202], Arize [253], Dynamo AI [55], Credo AI [212], Google [187], DataGrail [193], Dastra [225], OneTrust [224], Relyance AI [99], Transcend [100], Zendata [229], DeepMind SynthID watermark [45, 47]
Table 2: Overview of Research and Industry Practice in AI Safety Topics
5.3 AI Privacy
Privacy leakage in AI models refers to the unintended exposure of sensitive or private information that the models may have learned during training. Adversaries can exploit model weights or gradients to infer sensitive information in the training data, e.g., data reconstruction attacks, membership inference attacks, etc [288, 87, 281]. Also, as AI models might be trained on massive datasets that could contain personally identifiable information (PII) or confidential data, they might memorize specific details of personal information, which could be inadvertently reproduced during interactions with users. Moreover, malicious users can submit crafted adversarial prompts to the AI models to extract sensitive information, e.g., jailbreaking [117, 36].
Addressing privacy leakage in AI models involves implementing methods such as differential privacy [84, 236, 285, 284], federated learning [289, 282, 78, 280, 158], and privacy-preserving cryptographic protocols (e.g., homomorphic encryption [111, 180, 95]) for training AI models, ensuring that sensitive data cannot be extracted from weights or model outputs. Researchers also leverage differential privacy in finetuning or in-context learning [274, 283, 222], and design privacy-preserving prompts for querying LLMs [83, 52], to maintain data confidentiality during interactions between users and AI models. Other practices for ensuring data privacy include conducting rigorous data filtering before training, real-time monitoring of model outputs (i.e., moderation in §5.2), etc, such that the model outputs would not reveal sensitive information [133, 112, 80, 91] .
As we reflect on the practices of data privacy, a crucial aspect is the balance between privacy and utility. While ensuring data confidentiality is essential, it often comes at the cost of reduced functionality and utility. In practice, implementing privacy measures in AI services may not effectively prevent malicious users but inadvertently harm the experiences for benign users. This highlights a deeper issue: those determined to exploit vulnerabilities of AI services often find new, sophisticated ways to bypass safeguards, whereas being overly cautious may impact regular users negatively.
5.4 AI Security and Defense Methods
Defense methods typically involve the following strategies to ensure that AI models are robust and aligned with ethical guidelines: (1) employing prompt engineering methods, together with chain-of-thought reasoning or multi-agent systems, to help models better understand user inputs and reject harmful requests or unsafe outputs [176, 257, 247, 278]; (2) using built-in methods like exploring the decoding stage [88, 51], leveraging reward-based approaches [143, 188], or utilizing hidden states [29, 51] to prevent the generation of unsafe content; (3) constructing datasets with safety features, such as secure instructions or adversarial samples and training models with this data to improve their robustness against harmful inputs [262, 66, 133, 10, 168]; (4) aligning models with human values through Supervised Fine-Tuning (SFT) [287] and Reinforcement Learning from Human Feedback (RLHF) [157], which helps models to follow user instructions safely and adhere to desired ethical norms; (5) training models to detect unsafe content and developing guardrail pipelines to prevent models from taking unsafe inputs or generating inappropriate outputs (see §5.2); (6) employing certified robustness approaches that provide formal guarantees on the model’s robustness against adversarial perturbations [74, 40]; and (7) utilizing differential privacy  [1, 93, 209], homomorphic encryption [218, 95], secure MPC [44, 105], to ensure that sensitive data is protected during model training, etc.
5.5 AI Serving Security
AI serving security involves strategies to ensure secure deployment, operation, and maintenance of AI models when serving users [162, 258]. It focuses on protecting models and interactions from various threats, such as such as supply chain vulnerabilities [164], model theft [161], model denial of service [160], and insecure plugins [159].
Supply chain vulnerabilities exist in software components, models, and training data provided by third party providers, or user prompts collected with the supply chain [164]. Such vulnerabilities include using outdated models, deprecated third-party packages, or improper handling of training data, leading to data breaches, malicious injections, or even system failures [164]. Hackers may inject poisoned data through supply chain, introducing backdoors and biases during pretraining or fine-tuning [165]. Malicious users may submit malicious prompts to bypass system controls, inducing unauthorized actions or unsafe outputs of LLMs [163].
Model theft involves unauthorized access, replication, or reverse engineering of LLM weights and parameters to create functionally equivalent copies [161], posing risks to intellectual property, brand reputation, and financial security. Distributed denial of service (DDoS) attacks can compromise AI services, especially when models are integrated into widely accessible applications [160]. Attackers can craft prompts that exploit recursive behaviors in LLMs, which consumes excessive computational resources [190]. For example, an abnormal traffic pattern of a DDoS attack led to sporadic outages for OpenAI, which affected ChatGPT and developer tools for hours in November 2023 [166].
Plugins enable interactions with external software, databases, web tools, or APIs, which enhances LLM functionality [159]. However, they may introduce vulnerabilities if proper execution controls are lacking. Adversaries can exploit insecure plugins with adversarial prompts to execute unauthorized actions, such as data exfiltration, remote code execution, and privilege escalation [159]. Without proper input validation, plugins may reveal sensitive information [191]. Moreover, without isolated environments, unauthorized access to plugins could potentially modify system-level resources, which, in the worst case, might cause the underlying system to crash [191, 192].
5.6 AI Governance
AI governance is a system of rules, practices, processes, policies, and technological tools that ensure AI systems are built, developed, and used in a safe and responsible manner, aligning with social values, fulfilling legal standards, and meeting ethical principles [132, 21]. [214] defines 13 AI governance components in a continuous loop in AI life-cycle, including: 1) Establish Accountability for AI; 2) Assess Regulatory Risks; 3) Gather Inventory of Use Cases; 4) Increase Value of Underlying Data; 5) Assess Fairness and Accessibility; 6) Improve Reliability and Safety; 7) Heighten Transparency and Explainability; interpretability of AI. 8) Implement Accountability with Human-in-the-Loop; 9) Support Privacy and Retention; 10) Improve Security; 11) Implement AI Model Lifecycle and Registry; 12) Manage Risk; 13) Realize AI Value.
Among these directions, 1) and 8) are related to accountability; 2), 6), 9), 10), and 12) are related to jailbreaking and red teaming (§5.1), content moderation and guardrails (§5.2), AI privacy (§5.3), AI security and defense methods (§5.4), and AI serving security (§5.5); 3) is related to business and is out of the scope of this subsection (see [214] for details on practical AI governance use cases); 4) is related to data management such as training data processing [133], access control [203, 156, 16, 6], and data regulations [241, 14]; 5) is related to customizing AI workflows, i.e., guardrails (§5.2) and access control in AI systems [97, 6]; 7) is related to AI transparency; and 11) and 13) are related to customizing AI workflow for enhanced functionality and safety monitoring to the AI systems, i.e., guardrail (§5.2). Without loss of generality, this subsection discusses the rest topics, including i) AI accountability, ii) regulations, iii) access control, and iv) transparency, interpretability and explainability.
Accountability. Accountability refers to being responsible for actions and impacts of AI systems on individuals and society. It ensures responsibility and auditability, enables redress, and minimizes negative impact of AI systems [138]. However, defining and achieving accountability is challenging due to AI’s complexity and various considerations like privacy and intellectual property [187, 138]. [149] defines accountability as a relation of answerability with three conditions: authority recognition, interrogation, and limitation of power. It outlines a framework with seven features, including context, range, agent, forum, standards, process, and implications, as well as four key goals, including compliance, reporting, oversight, and enforcement. These goals are often complementary, while policy-makers tend to focus on some goals over others, depending on specific objectives of AI governance. Also, proactive approaches that set standards and anticipates issues, as well as reactive approaches that address and rectify issues after they occur, can be utilized to achieve the goal of accountability [149]. Accountability can be explored at different levels within the AI lifecycle, including data, model, and developers, and across various sectors such as government, private industry, and academia [186]. Approaches include algorithmic assessments, auditing, and data provenance techniques [186]. Watermarking methods further support accountability by embedding traceable identifiers in AI outputs, allowing responsible tracking of content in cases with ethical or regulatory implications [276, 277].
Regulations. AI regulation refers to laws, policies, and guidelines established to govern the development, deployment, and usage of AI technologies. European Union’s AI Act [238] categorizes AI applications by risk level, including unacceptable risk, high risk, general-purpose AI, limited risk, and minimal risk. High-risk AI applications, like those in healthcare and law enforcement, must meet high requirements of transparency, fairness, and safety. The General Data Protection Regulation (GDPR) [237] enforces strict data protection and privacy requirements. The Blueprint for an AI Bill of Rights [85] and NIST AI Risk Management Framework [150] provide guiding principles to foster responsible AI usage, emphasizing human oversight, transparency, and accountability.
Access control. Access control in AI systems manages user access to specific information within the model according to their permissions, thereby limiting unauthorized data exposure and ensuring compliance with data privacy regulations. Effective methods include data access controls [97] and model access controls [6]. As an example, compositional fine-tuning enables each information silo (e.g., databases or documents) has its own fine-tuned model, allowing models to operate securely across multiple data silos and ensuring users access only to the fine-tunings they are authorized for.
Interpretability, explainability, and transparency. Interpretability, explainability, and transparency are interrelated concepts but differ in their scope and focus. Interpretability focuses on understanding how an AI model functions internally and produces its outputs [124, 71, 194, 181]. Explainability provides explanations to specific AI outputs, presenting the results in ways that can be easily understood and trusted by end-users [37, 124, 71, 181, 264]. Transparency is the broadest concept that encompasses both interpretability and explainability and involves insights into the entire AI systems, such as how decisions are made, what data is used, and why specific results are produced [255, 110, 71, 181]. Transparency also extends to properties like traceability and content provenance that tracks models and data to ensure accountability and reliability [214, 45, 47]. In practice, these concepts are often considered together to provide a more holistic understanding of AI systems [291, 56, 71, 181, 233]. Practices to improve transparency, interpretability, and explainability include a variety of methods, including i) evaluating the contribution of various input elements (e.g., data features, user-defined concepts, or specific regions in images) to model outputs with methods such like removal-based explanations [41, 129, 124, 129, 201, 217, 46], counterfactual explanations [242, 98], concept activation explanations [103, 124], and saliency maps for visual data [3]; ii) X-of-Thought approaches that break down complex tasks into structured reasoning steps to make the reasoning process more interpretable to users, e.g., chain-of-thought [251, zhang2022automatic,diao2023active, 65], tree-of-thought [127, 271, 28], graph-of-thought [17, 113], and their variants; iii) embedding watermarks to enhance traceability and content provenance for AI-generated outputs [276, 277, 45, 47]; iv) leveraging external knowledge sources, such as retrieval-augmented generation (RAG) through vector databases [114, 94], function calling [62, 121], or web browsing [145], as reference documents for reasoning; v) leveraging interpretability algorithms or models as auxiliary models to generate insights into AI outputs and automate the interpretability of AI systems [195, 196, 20].
6 AI Safety in the 2020s: Challenges and Missions
This section utilizes the blueprint to guide our efforts today, navigating the challenges and opportunities of the current stage. As AI safety practitioners in the 2020s, we stand at a pivotal moment in history, with society moving toward an era where AI will be deeply integrated into daily life. While addressing current safety concerns, our primary mission is to anticipate and mitigate risks within the expanding AI landscape.
A shift from AI safety to AI quality assurance. The rapid changing AI landscape requires a shift of focus from narrowly defined safety concerns to a more comprehensive concept, i.e., AI quality assurance. Besides focusing on safety and security measures, practitioners can put more efforts on addressing alignment [157] of model outputs as well as the quality of interactions between humans and AI to enhance the overall quality of AI services. Viewing safety as part of the quality assurance concept helps to enhance AI in a way that is safe, accurate, and capable of meeting the complex demands of real-world AI applications. The terminology “quality assurance” also guide us to think more about what we can do to align our current efforts with the blueprint picture.
Alignment with more interdisciplinary collaborations. A key area that demands increased focus is the alignment of AI with deeper insights across different fields, people, and cultures. Current AI models, while advanced, often lack a deep understanding of the ethical, social, and cultural contexts, while most of the alignment research tend to be conducted by people in the computer science field. However, more advanced AI systems necessitate interdisciplinary collaborations between AI researchers and experts from different fields such as psychology, sociology, history, art, and anthropology, etc, especially for AI applications for specific fields such as finance [260, 282], healthcare [231, 2], education [184, 135], etc. AI safety practitioners must work towards bridging this gap by fostering interdisciplinary collaborations. By expanding our focus beyond technical aspects, we can develop AI systems that are more aligned with human values as well as the ethical, cultural, and social dynamics of our society.
Hallucination. Hallucination remains a fundamental and persistent challenge in AI systems, and is difficult to eliminate entirely, even as AI advances [89]. This is because hallucinations can be traced back to nearly every stage in the lifecycle of AI models, such as the quality and representativeness of training data, inappropriate user inputs, and the processes used during inference. Low-quality or biased training data can cause models to generate incorrect or misleading information, while inappropriate user inputs may lead to the creation of false details. During inference, models can also produce outputs that deviate from factual accuracy when they attempt to fill gaps in their training knowledge. Despite implementing Retrieval-Augmented Generation (RAG) [114] can mitigate hallucination, it remains an inherent risk of AI models, making it crucial for content moderation strategies to identify and address these errors consistently at the current stage. Achieving this requires an ongoing effort to refine training data, ground user queries properly, and implement moderation to minimize the negative impact of hallucinations on the quality of the model outputs.
Customization. Customization in AI workflow is essential for ensuring the AI systems to meet the diverse user needs while operating effectively in a variety of contexts. Current AI systems such as ChatGPT [133] tend to provide same services across all user groups, regardless of individual differences. However, as AI becomes more integrated into everyday life, one-size-fits-all solutions are no longer adequate. People have different backgrounds, preferences, and requirements, and AI systems should be able to adapt to individual differences and provide safe, effective, and relevant services. As an example, an AI model interacting with children should respond differently compared with interacting with adults, which requires to take into account age, learning objectives, cognitive abilities, emotional maturity, etc. Customization of AI services can enhance safety by adjusting contexts and interactions to align with specific cultural norms, legal regulations, and ethical considerations. In fields such as finance [260, 282], healthcare [231, 2], and education [184, 135], customizing AI services based on user profiles is crucial for preventing undesirable outputs and enhancing user experiences.
Interaction. From our long-term goal towards the blueprint future, the advanced society involves complex interactions between intelligent robots and human users (similar ideas are also mentioned in [11, 67]). This calls for a shift in our focus from static, one-shot analyses to dynamic, context-aware interactions between users and models to enhance alignment at the current stage. Besides prompt-based LLM systems or services, we can develop more sophisticated, context-aware interfaces for human-AI interactions to maintain coherent and multi-round interactions that adapts to evolving conversational contexts. Correspondingly, new safety protocols are in need to adjust in real-time to evolving contexts and user needs. It is also essential to develop new methodologies for safety assessment and mitigation that can function effectively in complex, rapidly evolving environments with multiple AI agents and human users.
7 Conclusion
This paper presents a blueprint for an advanced human society and leverages this forward-looking vision to guide today’s AI safety efforts. Through the blueprint, it becomes clear that Artificial General Intelligence (AGI) is not the ultimate goal of AI development. Instead, the true vision lies in the Internet of Everything, a deeply interconnected world where intelligent systems seamlessly integrate into every aspect of life. While AGI is a popular topic discussed among people in the computer science field today, a more advanced world demands interdisciplinary collaborations across various fields. Regarding AGI as the ultimate goal might limit our creativity in the 2020s, trapping us in “local optima” and potentially causing us to overlook the real challenges.
What will this future world be like? Will it bring more happiness to human-beings? The answer is uncertain. Advanced technologies may take on more duties of daily life, leaving humans with fewer tasks and more predictability in everyday life. On the other hand, with many everyday problems being solved by advanced technologies, people may have more time for self-reflection and personal growth, which may lead to another form of “happiness”.
As we stand at a pivotal moment in human history, our efforts become not only a technological endeavor but also a profound exploration of human potential and the boundaries of human capability. In any case, the journey toward this advanced future is fascinating, and every human being will look forward to it.

### v3

@@
 Abstract
-The growing prevalence of generative AI inevitably raises concerns regarding the associated risks and safety implications, which catalyzes significant progress in AI safety.
-However, as this field thrives, a critical question emerges: Are our current efforts aligned with the broader perspective of human history and civilization?
+The advancements in generative AI inevitably raise concerns about the associated risks and safety implications, which, in return, catalyzes significant progress in AI safety.
+However, as this field continues to evolve, a critical question arises: are our current efforts aligned with the long-term goal of human history and civilization?
 This paper presents a blueprint for an advanced human society and leverages this vision to guide contemporary AI safety efforts.
-It outlines a future where the Internet of Everything becomes reality, and create a roadmap of significant technological advancements towards this envisioned future.
+It outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future.
 For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face.
 By projecting current efforts against this blueprint, we examine the alignment between the present efforts and the long-term needs.
-This paper identifies gaps in current approaches and highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible future for AI development.
-This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby fostering AI’s role in promoting a more secure and sustainable future for human civilization.
+We also identify gaps in current approaches and highlight unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible and promising future of AI.
+This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a more secure and sustainable future in human civilization.
 1 Introduction
-The rapid development of AI and Large Language Models (LLMs) has fostered extensive progress in AI safety, specifically, LLM safety [171, 143, 51, 230, 90, 37].
-Researchers have been dedicated to addressing potential security and privacy risks in LLM lifecycle, aiming to align AI behaviors with human values and prevent misuse of LLMs, inappropriate outputs, information leakage, etc.
-However, despite the intense focus on LLM safety, a critical question emerges: Are current AI safety efforts truly aligned with the long-term evolution of human civilization, or are they simply addressing the concerns of today?
+The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety, specifically, LLM safety [175, 147, 54, 235, 93, 39].
+Researchers have been dedicated to addressing potential security and privacy risks in AI lifecycle, aiming to align AI behaviors with human values and prevent the misuse of AI models, inappropriate outputs, information leakage, etc.
+However, despite the intense focus on AI safety, a critical question emerges: are current AI safety efforts truly aligned with the long-term evolution of human civilization, or are they simply addressing the immediate concerns of the 2020s?
 One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [150].
-Despite their impressive capabilities in natural language processing and problem-solving [270, 265, 54, 38], today’s AI, including advanced LLMs [286, 260, 231], falls short of what could be considered genuine intelligence.
-These systems heavily rely on vast training datasets to function effectively, and lack true understanding, consciousness, self-awareness, and the ability to reason in ways comparable to human cognition.
-They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities capable of abstract thought and genuine reasoning.
-Moreover, recent studies have raised concerns about the reasoning abilities of LLMs, which may actually be a form of approximate retrieval from their training data and the deductive closure of the training data [96, 239].
-While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it fundamentally differs from human reasoning that involves abstract thinking, causal understanding, and the ability to generalize from limited examples.
-The unresolved issue of sustainable energy is another critical factor that challenges our current efforts to AI safety [5].
+Despite their impressive capabilities in natural language processing and problem-solving [280, 274, 57, 40], today’s AI, including advanced LLMs [297, 268, 236], falls short of what could be considered as “genuine intelligence”.
+Current AI models rely heavily on vast training datasets to function effectively, yet lack true understanding, consciousness, self-awareness, and reasoning abilities comparable to human cognition.
+They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities with authentic logical capabilities.
+Recent studies have raised concerns about the reasoning abilities of LLMs, which may be a form of approximate retrieval from their training data and the deductive closure of the training data [100, 245].
+While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs fundamentally from human reasoning that involves abstract thinking, causal understanding, and generalizing from limited examples.
+The energy issue is another critical factor that challenges our current efforts on AI safety [5].
 Current AI fails to fully represent a third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved.
@@
 However, current AI, rather than solving energy issues, consumes a significant amount of energy.
-Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent and approximately 0.02% of California’s annual electricity generation [7, 39, 169].
+Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent, and approximately 0.02% of California’s annual electricity generation [7, 41, 173].
 Also, inference with LLMs is computationally intense, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [144, 75].
 Given that ChatGPT alone has over 200 million weekly active users and receives over 1.54 billion page visits monthly [218], when scaled to billions of queries, the energy consumption becomes substantial.
-The significant energy requirements of current AI systems raise serious sustainability concerns and limiting their deployment throughout the human society.
-Until we overcome the energy constraints, AI’s ability to reshape society remains restricted.
+The significant energy consumption of current AI systems raises concerns about their sustainability, limiting their widespread deployment across human society.
+Until humanity overcomes the energy constraints, the potential of AI to reshape society will remain limited and uncertain.
 Figure 1: Illustration of Internet of Everything in a Household
-These concerns suggest that our efforts on AI safety today may fail to capture deeper insights that could emerge as AI evolves, potentially neglecting technological challenges that ultimately shape the impact of AI on human civilization.
+These concerns indicate that our current efforts in AI safety might overlook more complex issues that could arise as AI continues to evolve.
+We potentially fail to capture deeper insights that will ultimately shape the impact of AI on human civilization.
 This raises several critical questions:
-  • Are we focusing on the right problems that align with the long-term AI development?
+  • Are we focusing on the right problems that align with the long-term advancements of AI?
   • Are we potentially overlooking significant challenges that will emerge as AI evolves?
-  • Are we on the right track for addressing real challenges in AI safety?
+  • Are we on the right track of addressing real challenges in AI safety?
 “The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety research.
-In this advanced future, energy issues have been resolved, and intelligent chips and brain-computer interfaces have been fully developed, which enables advanced intelligent robots to integrate into everyday life.
-Breakthroughs in foundational theories, such as Einstein’s relativity [57, 58] and Quantum Mechanism [59, 279], have catalyzed revolutionary progress in AI, and humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1.
-While this vision may be beyond the horizon of several decades, by reviewing the present landscape of AI safety against this far future, we can gain deeper insights into the potential and limitations of our current approaches, as well as specific missions for AI safety researchers in the 2020s.
-It reminds us that, while addressing today’s immediate concerns, we should avoid being constrained by these short-term solutions and look toward the future to guide our present efforts.
+In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity [60, 61] and Quantum Mechanism [62, 289], have catalyzed revolutionary progress in AI.
+Intelligent chips [89, 119] and brain-computer interfaces [84, 151, 262] have been fully developed, and advanced intelligent robots are integrated into everyday life.
+Humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1.
+While this vision may be beyond the horizon of several decades, by reviewing the present landscape of AI safety against this far future, we can gain deeper insights into the potential and limitations of our current efforts, as well as specific missions for AI safety researchers in the 2020s.
+It reminds us that, while addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, look toward the future to guide our present efforts.
 Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” to encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape.
-ii) Deeper collaboration with experts from diverse fields is essential to align AI with human values better.
-iii) AI workflow customization and user-AI interactions are important, especially in dynamic contexts and across different user groups.
+ii) Deeper collaborations between experts from different fields are essential to align AI with human values better.
+iii) AI and AI safety workflow customizations are crucial to ensure that AI systems can adapt to diverse user requirements while maintaining safety and reliability across different use cases.
+iv) Exploring and understanding interactions between users and AI are crucial for enhancing user experiences, especially in dynamic contexts across different user groups.
 Humanity’s pursuit of higher technology will never stop. While searching for the sixpence on the ground, do not forget to look up at the night sky and look for the moon.
@@
 2 Related Works
-Multiple AI safety surveys and benchmarks have been published in recent years.
-These studies focus on different aspects of AI safety, such as jailbreaking, red teaming, guardrail, ethics, etc.
+Multiple AI safety surveys and benchmarks have been published in recent years [106, 8, 126, 39, 189, 128, 85, 282].
+These studies focus on different aspects of AI safety, such as jailbreaking, red teaming, guardrail, ethics, etc [128, 282, 146].
 A summary of these works is presented in Table 1, which provides an overview of the related works and their focus within the field of AI safety.
@@
 [106] provides valuable enterprise insights into the practical challenges and opportunities for AI safety in industries.
-[67] provides a systematic analysis of the ethical risks of advanced AI assistants, and discusses their potential influence on users, the risks of misuse, strategies for enhancing human-AI interactions, and the broader implications for human society.
-[43] proposes the concept of “Guaranteed Safe” (GS) AI, emphasizing the need for formal, verifiable, and auditable safety guarantees to ensure robust AI behavior.
+[70] provides a systematic analysis of the ethical risks of advanced AI assistants, and discusses their potential influence on users, risks of misuse, strategies for enhancing human-AI interactions, and the broader implications for human society.
+[45] proposes the concept of “Guaranteed Safe” AI, emphasizing the need for formal, verifiable, and auditable safety guarantees to ensure robust AI behavior.
 [11] proposes a technical roadmap towards next-generation human-centered AI safety that aims to anticipate and prevent potential hazards in the interactions between AI systems and users.
-Specifically, the paper bridges the fields of AI safety and control systems, and highlights the open challenges at their intersection.
+This paper bridges the fields of AI safety and control systems, and highlights the open challenges at their intersection.
 The authors emphasize the importance of leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for achieving meaningful safety assurances.
-Different from existing works, this paper envisions an advanced society in the future of human civilization, and use this vision to guide current AI safety research.
-By comparing the current landscape of AI safety with this distant future, this work identifies both the potential and limitations of today’s approaches and suggests specific missions for AI safety researchers in the 2020s.
-This perspective encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
+Different from existing works, this paper envisions an advanced society in the future of human civilization, and use this vision to guide current AI safety efforts.
+By aligning the current landscape of AI safety with this distant future, we identify the potential and limitations of today’s efforts, and suggest specific missions of AI safety researchers in the 2020s.
+This paper encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
 3 From the New World
-In this future, humanity has entered an unprecedented era.
-Scientists have achieved breakthroughs in energy production with technological innovations.
-The resulting flood of abundant, clean energy has revolutionized human civilization beyond the our ancestors’ wildest dreams, reshaping the world and initiating a new chapter of human civilization.
+In this future, humanity has entered an unprecedented era, driven by technological breakthroughs in energy generation.
+The flood of clean energy has revolutionized human society beyond the wildest dreams of our ancestors, reshaping the world and initiating a new chapter of human history.
 3.1 Imagine the Future: A Blueprint
-The transformative energy revolution has catalyzed human civilization from every aspect.
-Foundational theories of science have undergone revolutionary breakthroughs, with major advancements in nanotechnology [136, 19, 61], Quantum Mechanics [59, 279], and Einstein’s relativity theories [57, 58].
-The world has been reshaped by innovative technologies such as intelligent chips [86, 115], brain-computer interfaces [147, 81, 254], holographic technology [22, 273, 15], and advanced 3D printing [205, 73], bringing changes in daily life of miraculous.
+The energy revolution has transformed nearly every aspect of human society.
+Foundational theories have experienced groundbreaking advancements, particularly in fields such as nanotechnology [140, 19, 64], quantum mechanics [62, 289], and Einstein’s relativity theories [60, 61].
+Innovative technologies, such as intelligent chips [89, 119], brain-computer interfaces [151, 84, 262], holographic technology [22, 283, 15], and advanced 3D printing [210, 76], have reshaped the world, bringing about profound changes in daily life that were once beyond our imagination.
 Transportation has evolved dramatically.
-While simple mechanical vehicles, such as fixed-track trains and high-speed rail [144, 179, 109], still exist due to their efficient and straightforward design, they operate at astonishing speeds with high stability, powered by efficient and limitless energy.
-Advanced autonomous vehicles and aircraft are widespread and are controlled by intelligent chips, eliminating the need for any human intervention.
-Also, due to breakthroughs in Einstein’s relativity theories [57, 58] and Quantum Mechanics [59, 279], humanity has unlocked the potential of instant transportation that allow for long-distance travel in short time that transcends the imagination of the 21st century.
-Advanced brain-computer interfaces [147, 81, 254] and intelligent microchips [86, 115] has revolutionized the way people perceive, which expands the boundaries of human cognition.
-No longer are people operating devices and asking questions in search engines or LLMs; the embedded intelligence chips help them to gather more comprehensive information and interact with the surrounding environment.
-Meanwhile, communication over long distances has transcended the limitations of cables and the outdated internet; information, thoughts, emotions, and sensations flow through a new medium due to the advancement in quantum computing [77, 198].
-Intelligent robots have been integrated into every aspect of modern society and play important roles of daily life; see Figure 1.
-Different from their “data-hungry machine intelligent ancestors” that relies on huge amount of training data and statistical patterns in the early 21st century [131, 286], these intelligent robots perceive and learn through intelligent chips, which are real artificial intelligence and possess genuine understanding and adaptability, much like a human child discovering the world for the first time.
-Also, their functionalities are customizable through specialized chips based on user needs.
-Robots can engage in complex interactions with people and adapt their capabilities and expertise to suit their human counterparts’ requirements.
-In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds have been blurred, i.e., Internet of Everything.
-Such a network integrates objects with embedded intelligent chips, such as smart devices, autonomous systems, intelligent robots, and humans cognition, which allows for unprecedented levels of efficiency, personalization, and adaptability and fundamentally reshapes the way people interact with the world.
+While simple mechanical vehicles, such as fixed-track trains and high-speed rails [148, 183, 113], still exist due to their efficient and straightforward design, they operate at astonishing speeds with high stability, powered by abundant efficient energy.
+Advanced autonomous vehicles and aircraft, controlled by intelligent chips, are widespread, eliminating the need for human intervention.
+Also, with breakthroughs in Einstein’s relativity theories [60, 61] and quantum mechanics [62, 289], humanity has unlocked the potential of instant transportation that allow for long-distance travel in short time that transcends the imagination of the 21st century.
+Advancements in brain-computer interfaces [84, 151, 262] and intelligent microchips [89, 119] have revolutionized the way people perceive and expanded the boundaries of human cognition.
+No longer are people operating devices and asking questions in search engines or AI models such as LLMs; the embedded intelligence chips help people to gather comprehensive information and interact with the surroundings.
+Also, communication over long distances has transcended the limitations of the outdated cables and the internet; information, thoughts, emotions, and sensations flow through a new medium due to the advancements in quantum computing [80, 203].
+Intelligent robots have been integrated deeply into human society and play important roles in daily life; see Figure 1.
+Different from their “data-hungry machine intelligent ancestors” that rely on huge amount of training data and statistical patterns in the early 21st century [135, 297], these intelligent robots perceive and learn through intelligent chips.
+They are real artificial intelligence that possess genuine abilities of learning, understanding, and adapting, much like a human child discovering the world for the first time.
+Their functionalities are customizable through specialized chips based on user needs, and they can engage in complex interactions with people and adapt their capabilities and expertise according to their human counterparts’ requirements.
+In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds are blurred, i.e., the Internet of Everything.
+Such networks integrate objects embedded with intelligent chips, e.g., smart devices, autonomous systems, intelligent robots, and humans, allowing for unprecedented levels of efficiency and adaptability and fundamentally transforming the way people interact with the world.
 3.2 Retrospective on AI Safety in the New World
-In a world where the Internet of Everything has become a reality, intelligent robots are customized to meet diverse user needs and seamlessly integrate into daily life; see Figure 1.
-Consequently, the concept of AI safety has been redefined to focus more on the quality of services they provide, i.e., emphasizing comprehensive quality assurance.
-Unlike today’s command-and-response interactions between humans and LLMs, future robots can engage in sophisticated reasoning, emotional cognition, and possess appropriate levels of autonomy.
-They will also interact dynamically with other entities, including humans and fellow robots, within the interconnected system.
-The Internet of Everything scenarios necessitate the robots to handle complex interactions, demanding a sophisticated understanding of the environment for decision-making.
+In a world where the Internet of Everything has become a reality, intelligent robots are customized to meet diverse user needs and are seamlessly integrated into daily life; see Figure 1.
+Consequently, the concept of AI safety has been redefined to prioritize the quality of services provided by the intelligent robots, i.e., emphasizing comprehensive quality assurance rather than solely focusing on safety, security, and privacy risks.
+Unlike today’s command-and-response interactions between humans and AI models, future robots can engage in sophisticated reasoning and emotional cognition, and possess appropriate levels of autonomy.
+They will also interact dynamically with other entities, including humans and fellow robots, within the interconnected network.
+The Internet of Everything scenarios necessitate the robots to handle complex interactions, thus requiring a sophisticated understanding of the surroundings for decision-making.
 Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level.
-Such advanced interactions require deep understanding of the environments, demanding real-time information retrieval and processing, which is far beyond what can be achieved with simple code logic.
-Robots must continuously collect and analyze data from their environment, and integrate external inputs with contextual knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align precisely with human needs and expectations.
+Such advanced interactions requires the intelligent robots to retrieve and process information in real-time, which is far beyond what can be achieved with simple code logic.
+Robots must continuously collect and analyze data from the surroundings and integrate external inputs with contextual knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align with human expectations.
 Customization is a key aspect of future robots for providing personalized services to meet individuals’ requirements.
-Even robots of the same type should be capable of adapting their services to different individuals.
-Teaching robots should adjust educational materials and teaching methods to capitalize on each student’s unique talents and strengths, while cleaning robots should intelligently schedule their works to minimize disruption to other humans in the household, e.g., avoiding occupied workspace.
-Such customization aligns the robots’ capabilities with specific user preferences, and ensures that the robots serve as effective assistants across a wide range of daily tasks and environments.
+Even robots of the same type should be able to adapt their services to different individuals.
+Teaching robots should adapt educational materials and teaching methods to leverage each student’s unique talents and strengths, while cleaning robots should intelligently schedule their tasks to minimize disruption to household members, such as avoiding occupied workspace.
+Such customization aligns the robots’ capabilities with specific user needs, ensuring that the robots serve as effective assistants across a wide range of daily tasks and complex environments.
 Finally, the effectiveness of intelligent robots requires a balance between robot autonomy with human oversights.
 Too little autonomy would burden users with continuous supervision, while too much autonomy could raise safety concerns or result in unintended behaviors.
-For instance, a care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions.
-Conversely, while the tasks of cleaning robots are generally routine, they must still adapt to specific circumstances, such as recognizing when an area is occupied by people and postponing cleaning activities.
-An appropriate level of robot autonomy optimizes the division of labor between humans and robots, ensuring that robots function effectively and safely while enhancing our daily lives without compromising human control over these intelligent robots.
+A care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions.
+Conversely, while the tasks for cleaning robots are typically routine, they still need to adapt to specific situations, such as detecting when an area is occupied and postponing cleaning activities accordingly.
+An appropriate level of robot autonomy ensures an effective division of labor between humans and robots, allowing robots to operate efficiently and safely while enhancing our daily life without compromising human control over these intelligences.
 4 Bridging Today and the Future
 Figure 2: Evolution of Human Civilization
-This section presents an ambitious roadmap towards the aspirations for the advanced future, exploring key milestones in the evolution of AI and human civilization.
-By considering potential paradigm shifts in fundamental theories and breakthroughs in technology, we feature this fascinating journey into three eras, i.e., i) the data-driven era where AI learns knowledge from vast training datasets, ii) the neural augmentation era where human-machine interface [147, 81, 254] and intelligent chips [86, 115] extend the boundary of human cognition, and iii) the revolutionary era with a breakthrough in energy production that catalyzes transformative changes across every aspect of human society; see §4.1, §4.2, and §4.3, respectively.
-We then examine the evolving paradigms of AI safety throughout this journey, analyzing the potential focuses and challenges of each era and exploring the impacts of technological advancement on AI safety in §$4.$
+This section outlines an ambitious roadmap to the advanced future, highlighting key milestones in the evolution of AI and human civilization.
+By exploring potential paradigm shifts in fundamental theories and technological breakthroughs, this section categorizes this journey into three eras, i.e., i) the data-driven era, in which AI learns from vast training datasets, ii) the neural augmentation era, where human-machine interface [151, 84, 262] and intelligent chips [89, 119] expand the boundaries of human cognition, and iii) the revolutionary era, characterized by a breakthrough in energy production that catalyzes transformative changes across all aspects of human society; see §4.1, §4.2, and §4.3, respectively.
+We then examine the evolving paradigms of AI safety throughout this journey, analyzing the focuses and challenges of each era while exploring the impacts of technological advancement on AI safety in §4.4.
 4.1 Rookie Level: Data-Driven Era
-The Data-Driven Era represents the early phase of AI development, where AI models heavily rely on the quality, availability, and quantity of training data [131].
+The data-driven era represents the earliest phase of AI development, where AI models rely heavily on the quality, availability, and quantity of training data [135, 261, 269].
 AI makes probabilistic inferences based on patterns found in massive training datasets, thus failing to achieve genuine understanding of data [100, 245].
-While AI in this era cannot be regarded as real intelligence, it assists people to solve problems and enhances productivity in the society.
+While AI in this stage cannot be regarded as real intelligence, it can still assist people to solve problems and enhances productivity in human society.
 Stage 1: Past, Current, and the Near Future.
-This stage is characterized by diverse AI models for processing various forms of data across different scenarios [131, 50, 63, 210, 231, 260].
-Researchers have been engaging in refining model structures and developing advanced algorithms, enabling more sophisticated analysis of data in different formats.
-The thrive of LLMs exemplifies this progress, showcasing unprecedented capabilities of AI models in language processing and generation [260, 286, 231].
-Concurrently, we observe significant advancements in other fields such as computer vision [101, 221, 256], with progress in image processing [235, 25, 216] and video analysis [223, 261, 267], as well as advancements in audio recognition [64, 178, 4, 106].
-This stage is driven by exploring vast datasets and the refining algorithms continuously, which consistently pushes the boundaries of AI capabilities within specific domains.
+This stage is characterized by developing diverse AI models to process various forms of data across different scenarios [135, 53, 66, 215, 236, 268].
+Researchers have been engaging in refining model architectures and developing advanced algorithms, enabling more sophisticated analysis of data in different formats.
+The thrive of LLMs exemplifies this progress, demonstrating unprecedented capabilities of AI models in language processing and generation [268, 297, 236].
+Concurrently, significant advancements have also been observed in other domains such as computer vision [105, 226, 264], image processing [241, 25, 221], video analysis [228, 270, 276], and audio recognition [67, 182, 4, 110].
+This stage is driven by the exploration of vast datasets and the continuous refinement of algorithms, which consistently pushes the boundaries of AI capabilities across specific domains.
 Stage 2: Interdisciplinary AI.
-This stage is characterized by leveraging AI to transform traditional fields, such as healthcare [231, 2], finance [260, 282], and education [184, 135], through cross-disciplinary applications in human society.
-Mature multi-agent systems [116, 246, 263, 79] that leverage AI entities with different specializations will be widely utilized to assist people and solve sophisticated problems.
-Such agents participate in decision-making processes, or represent users to finish some tasks, providing more sophisticated solutions of a given problem.
-In healthcare, for example, a multi-agent system may involve collaboration between various medical agent specialties and integrate data from diverse sources to formulate comprehensive treatment plans.
-The rise of interdisciplinary AI integration will amplify the utility of AI in more complex scenarios.
-Although AI at this stage is still not real intelligence, it becomes deeply integrated in different domains in human society, and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
+This stage is characterized by the integration of AI through cross-disciplinary applications, driving significant innovations and transformations in traditional fields such as healthcare [236, 2], finance [268, 292], and education [188, 139].
+In this stage, AI applications would unlock new opportunities across industries and reshape human society.
+Mature multi-agent systems [120, 253, 272, 82] that leverage AI entities with specialized capabilities might be widely deployed to assist people to solve sophisticated problems.
+Such agents will participate in decision-making processes or represent users to complete some tasks.
+For example, in healthcare, a multi-agent system may involve collaborations between multiple specialized medical agents and integrate data from various sources to formulate comprehensive treatment plans.
+The rise of interdisciplinary AI integration will enhance AI’s utility in increasingly complex scenarios.
+While AI at this stage does not yet achieve true intelligence, it becomes deeply integrated across various domains in human society and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
 4.2 Champion Level: Neural Augmentation Era
-In this era, AI models have been widely applied in human society, and humans may witness breakthroughs in advanced technologies and fundamental theories, which may extend the boundary of human cognition and life experience.
+In the neural augmentation era, humans may witness breakthroughs in fundamental theories and advanced technologies, which may extend the boundary of human cognition and enhance life experience.
 Stage 3: Breakthroughs in Advanced Technologies.
-In this stage, breakthroughs in advanced technologies emerge, such as brain-computer interfaces [147, 81, 254], nanotechnology [136, 19, 61], and holographic technology [22, 273, 15].
-These innovations enable unprecedented levels of interaction between humans and machines by augmenting human senses and cognition.
-For instance, intelligent chips [86, 115] may be developed that integrate directly with the human body, creating a seamless interface between biological and digital systems.
-Such advances also change the nature of the data.
-Unlike traditional static formats such as text, images, or videos, the data in this stage becomes more comprehensive, encompassing dynamic, real-time, and sensory-rich formats, such as neural activity patterns or real-time environmental data from augmented reality systems [31, 32].
-In addition to the advancements in AI seen during the Data-Driven Era, this stage further expands the boundaries of human perception and cognition, revolutionizing how people experience and interact with the world.
-However, widespread application of these techniques is limited by energy constraints and the significant resources required for implementation.
+At this stage, breakthroughs in advanced technologies, such as brain-computer interfaces [84, 151, 262], nanotechnology [140, 19, 64], and holographic technology [22, 283, 15], augment human senses and cognition, enabling unprecedented levels of interactions between humans and machines.
+Language barriers may disappear, as human can perceive with chips for understanding, which enables ubiquitous real-time context-aware translations between people of different languages and cultures.
+Intelligent chips [89, 119] might be integrated into human bodies, creating a seamless interface between biological and digital systems.
+Such advancements may change the nature of the data.
+Unlike data of static formats such as text, images, or videos, the data might be more comprehensive, in dynamic and sensory-rich formats, such as neural activity patterns or real-time environmental data from augmented reality systems [31, 32].
+This stage expands the boundaries of human perception and cognition, revolutionizing how people experience and interact with the world.
+However, applying these techniques will be still challenging due to the energy constraints and the significant resources required by these technologies.
 Stage 4: Breakthroughs in Fundamental Theories.
-In this stage, humanity experiences a new wave of breakthroughs in fundamental theories, such as Einstein’s Relativity Theories [57, 58] and Quantum Mechanics [59, 279].
-These theoretical advances redefine our understanding of the universe and offer new possibilities for innovations in technologies.
-Such breakthroughs may also unlock solutions to longstanding challenges in human history, i.e., the energy issues, offering clean, abundant, and sustainable energy sources.
+At this stage, humanity may witness a new wave of breakthroughs in fundamental theories, such as Einstein’s theories of relativity [60, 61] and quantum mechanics [62, 289].
+These theoretical advancements have the potential to redefine our understanding of the universe, and, moreover, offer new possibilities for groundbreaking technological innovations and provide critical insights into solving the longstanding challenges in human history, i.e., the energy issues.
 4.3 Ultimate Level: Revolutionary Era
 Stage 5: Energy Revolution.
-A new era arrives as humanity finally achieves energy revolution, unraveling the mysteries of sustainable power generation.
+A new era would arrive as humanity finally achieves energy revolution, unraveling the mysteries of sustainable power generation.
 This may be achieved through mature technologies that are currently on the horizon, such as breakthroughs in large-scale controlled nuclear fusion [73, 205, 187], or revolutionary solar technologies [23, 141], together with advancement in nanotechnologies [140, 19, 64] that produce new materials for collecting and storing solar energy efficiently.
@@
 These satellites might then transmit the abundant energy to the receivers on Earth.
-Regardless of the specific technology, the flood of clean, abundant energy can eliminate the dependence on fossil fuels, and fundamentally revolutionize the global economic landscapes as well as human civilization.
+Regardless of the specific technology, the flood of clean, abundant energy can eliminate our dependence on fossil fuels, and revolutionize the global economic landscapes as well as human civilization.
 Stage 6: The Internet of Everything.
-The final stage is the realization of Internet of Everything.
-Every object, living being, and system becomes part of a vast, intelligent network.
-This hyper-connected world is characterized by unprecedented levels of perception, communication, and interactions between humans and intelligent robots in an area.
-Language barriers may disappear, as human can perceive with chips for understanding, which enables ubiquitous real-time context-aware translations between people of different languages and cultures.
-Due to everyone is connected to a huge network, the boundary of human sense has been extended, which may also catalyze breakthroughs in holographic technology [15, 22, 273].
+The final stage is the Internet of Everything.
+Every object, living being, and system become parts of a vast, intelligent network.
+This hyper-connected world is characterized by unprecedented levels of perception, communication, and interactions between humans and intelligent robots.
+In this era, everyone is connected to a huge network, and the boundary of human sense has been extended, which may also catalyze breakthroughs in holographic technology [15, 22, 283].
 Remote experiences become true; working, entertainment, and communication are revolutionized, and human has entered the ultimate advanced era.
@@
 Current AI safety practices primarily focus on data safety and model safety [39, 106, 44, 225, 171, 198, 260, 230, 229].
-A critical aspect involves data privacy and security, particularly given the vast amounts of sensitive information used in AI training [118, 120, 83, 117].
-AI safety researchers and practitioners also explore adversarial attacks to training data or AI models [9, 35, 119, 244, 211, 269], as well as malicious inputs designed to manipulate AI systems or induce producing unsafe content, e.g., jailbreaking or prompt injection attacks [75, 292, 126, 33, 90, 258, 293].
-Correspondingly, red-teaming ML models [12, 68, 173, 197, 243] is widely used to identify potential vulnerabilities in ML models with attacks to further enhance the robustness of AI models, and significant work has been proposed to enhance security and robustness of AI systems [27, 240, 290].
-Additionally, researchers have been engaging in AI governance to enhance AI accountability [21], transparency and explainability [291] in AI systems to ensure the decision-making processes are understandable, accountable, and aligned with regulations.
-Other significant areas of focus is aligning AI outputs with human values [249, 174, 232, 206, 245].
+AI safety researchers and practitioners have explored adversarial attacks targeting training data and AI models [9, 36, 123, 250, 216, 279] to achieve malicious goals, for example, designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection attacks [78, 303, 130, 34, 93, 266, 304].
+Correspondingly, red-teaming ML models [12, 71, 177, 202, 249] is widely used to identify potential vulnerabilities in AI models, and significant work has been proposed to enhance security and robustness of AI systems [27, 246, 301], aligning AI with human values [256, 178, 237, 211, 252].
+Data privacy is also a critical aspect, given the vast amounts of sensitive information in the training data [122, 124, 86, 121].
+Additionally, researchers have been engaging in AI governance to enhance AI accountability [21], transparency and explainability [302] in AI systems to ensure the behaviors of AI are understandable, accountable, and aligned with regulations [21, 136, 219].
+.
 Neural Augmentation Era.
-Advancements in cognitive chips and brain-computer interfaces have extended the boundaries of human cognition and have changed the way of human perceiving the world.
-With the integration of intelligent chips and their direct neural connections to AI systems, users can interact with AI with high flexibility, and the customization of AI services towards different user groups or domain knowledge might be a critical focus [248].
-Individuals might rely on AI in different ways based on their background and cognitive needs, thus, personalized experiences of AI services should be explored for different user needs and preferences.
+Cross-disciplinary AI applications might have been widely deployed in human society, and advancements in intelligent chips [89, 119] and brain-computer interfaces [84, 151, 262] might extend the boundaries of human cognition and have changed the way of human perceiving the world.
+With the integration of intelligent chips and their neural connections to intelligent systems, users can interact with machines and AI with high flexibility, and the customization of AI services for different user needs might be a critical focus [255].
+Individuals might rely on AI in different ways, and personalized experiences of AI services might be explored based on users’ background and requirements.
 Revolutionary Era.
-In the revolutionary era, AI safety challenges may focus on the complex interactions between interconnected entities, such as humans and intelligent robots, within the network.
-As users may have diverse needs and robots will provide customized services, ensuring the safety and reliability of these interactions becomes challenging.
-Robots may need to request and integrate data from multiple external sources in real-time to fulfill user requests.
+In the revolutionary era, AI safety challenges may focus on complex interactions between interconnected entities, e.g., humans and intelligent robots, within the network.
+Robots might provide customized services according to different user requirements, and ensuring the safety and reliability of the interactions between them becomes challenging.
+Robots may need to integrate data from external sources in real-time to fulfill user requests.
 Managing the quality and security of these interactions is crucial, as any breakdown or interference could lead to incorrect or even harmful outcomes.
-Ensuring seamless communication, data integrity, and decision-making in such a sophisticated network may require smarter, more efficient methods, such as advanced cache management systems, to support the dynamic and complex interactions.
-This requires addressing issues of latency, data accuracy, and the security of transmitted information to maintain a safe and reliable network to ensure high-quality interactions between humans and robots in the network.
-By aligning AI with the progress of human civilization, we can see the path of AI safety evolves in different stages, which reflects the growing complexity of the roles of AI in human society.
+Thus, ensuring seamless communication, data integrity, and efficient decision-making in such a sophisticated network may require smarter and more efficient methods for managing, such as advanced cache management systems, to support the dynamic and complex interactions between humans and robots in the network.
+By aligning advancements of AI with human civilization, we can see the focuses of AI safety evolve in different stages.
 Initially, the focus is on addressing safety issues related to data and models.
-As AI and novel technologies become more developed, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs.
-Finally, when entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the same network.
-These insights offer guidelines for what we can do for AI safety at the current stage.
-In the following sections, we will overview today’s AI safety landscape, then reflect on the progress in this area so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed in the present stage.
+With progress of AI and advanced technologies, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs.
+Finally, when humanity entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the network.
+These insights offer guidelines for our efforts on AI safety at the current stage.
+The following sections will overview today’s AI safety landscape, then reflect on the progress in this area so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed at the current stage.
 5 Back to the 2020s: Today’s AI Safety Landscape
-This section overviews mainstream AI safety practices so far, and project them against the blueprint future in §3.1 to guide our current efforts in AI safety.
+This section overviews current mainstream AI safety practices, and project them against the blueprint future in §3.1 to guide our current efforts on AI safety.
 5.1 Jailbreaking and Red Teaming
 Jailbreaking [51, 127, 179, 212, 220, 224, 240, 257, 266, 277, 304] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs.
-Malicious users or adversaries utilize these techniques to manipulate the model into producing abnormal or harmful responses, such as giving instructions for illegal activities and generating offensive content.
-Common jailbreak methods include crafting inputs to trick the model, planting backdoors to exploiting vulnerabilities of the model, etc.
-Such activities pose significant risks by undermining safety measures of AI models, making it challenging to ensure the model to behaves appropriately across all scenarios.
-Red-teaming is a common practice for mitigating unsafe behaviors in LLMs, where teams of experts simulate attacks or misuse scenarios to identify potential vulnerabilities in AI models [68, 243, 197, 12, 173, 128, 275].
-By actively challenging AI models and probing for weaknesses through adversarial inputs, bias exploitation, or system manipulation, it helps uncover potential risks that may not be evident in normal usage, such as biases, or failure modes that might lead to undesirable behaviors in AI models.
-The outputs from red-teaming can be used to train the model to be less likely to cause harm or steer it away from undesirable outputs.
-This proactive approach has proven effective in stress-testing AI models and has been widely used in industry practices [18, 173, 68].
-As AI systems grow more sophisticated, red teaming must continually adapt to keep pace with increasingly complex attacks, which calls for more advanced tools and interdisciplinary expertise [153].
-Also, model owners can conduct regular and diverse red teaming to anticipate and mitigate risks in evolving use cases.
-Moreover, current red-teaming practices are often static, typically conducted after model pretraining and consisting of one-time attacks.
-While this approach provides useful insights, it does not fully account for the evolving user needs, especially when dealing with customized AI services.
-To address this, red teaming should simulate diverse user groups and scenarios, such as simulating a diverse range of user groups and their unique needs, to address the customization of AI services and enable a comprehensive evaluation of potential vulnerabilities across different contexts.
-5.2 Content Moderation and Guardrails
-Content moderation involves detecting and reviewing user input and AI-generated content to ensure it complies with safety, ethical, and regulatory standards [133, 112, 80, 91].
-Moderation approaches leverages rule-based methods, machine learning classifiers, and human oversight to monitor, evaluate, and manage inputs and outputs of AI models.
-These approaches identify and address inappropriate content, such as toxicity, bias, hallucinations, sensitive information, and jailbreaks, to ensure that the AI models do not produce or engage with harmful content, which maintains responsible AI behaviors.
-Guardrails [141, 189] employ customizable AI workflows for ensuring the quality of the model outputs while enhancing the control of the whole workflows.
-It functions as an intermediary layer between users and AI models, enabling users to add components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
-As we reflect on practices for AI moderation and guardrails, several key challenges and considerations have emerged.
-The balance between utility and safety is crucial and may require dynamic solutions according to the changing contexts.
+Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content.
+These activities pose significant risks by compromising safety measures of AI models, making it challenging to ensure the model to behave appropriately, especially in real-world scenarios where the potential inputs are unpredictable.
+Red-teaming is a common practice that simulates attacks or misuse scenarios to identify vulnerabilities in AI models, thereby mitigating unsafe behaviors [71, 249, 202, 12, 177, 132, 285].
+It challenges AI models with adversarial inputs, bias exploitation methods, and system manipulation to uncover potential risks, e.g., hidden biases or failure modes, that may not be evident during normal usage.
+The outcomes from red-teaming can be leveraged to refine AI models, reducing undesired behaviors and inappropriate outputs.
+This approach is effective in stress-testing AI models and has been widely employed in industry [18, 177, 71].
+As AI systems grow more sophisticated, red-teaming must continually adapt to counter increasingly complex attacks, which calls for more advanced tools and interdisciplinary expertise [157].
+Model owners can conduct diverse red teaming regularly to anticipate and mitigate risks in evolving use cases.
+Also, current red-teaming practices are often static, typically conducted after model pretraining with one-time attacks.
+Thus, these practices potentially fail to address evolving user needs and complex interactions between users and models, especially when dealing with customized AI services for different user groups.
+To address this, red teaming should involve more diverse use cases, such as employing different user groups and simulating complex interactions with dynamic contexts, to address safety risks in customized AI services and enable a comprehensive evaluation of potential vulnerabilities across different scenarios.
+5.2 Content Moderation and Customized Moderation Pipeline
+Content moderation involves detecting and reviewing both user input and AI-generated content to ensure compliance with safety, ethical, and regulatory standards [137, 116, 83, 94].
+Moderation approaches employ rule-based methods, machine learning classifiers, and human oversight to monitor, evaluate, and manage the inputs and outputs of AI models.
+These approaches identify inappropriate content, such as toxicity, bias, hallucinations, private information, and jailbreaks [251].
+By preventing AI models from engaging with or producing harmful content, content moderation plays a critical role in maintaining responsible AI behavior.
+The customized moderation pipeline [145, 194] employs multiple components at different stages of LLM inference to enhance the quality of the model outputs while enhancing the control of the whole AI workflows.
+It functions as a flexible intermediate layer between users and AI models, enabling users to add customized components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
+As we reflect on practices for AI moderation and customized moderation pipelines, several key challenges and considerations have emerged.
+1) Hallucination remains a significant issue and is unlikely to solve completely, regardless of how AI evolves.
+That is because hallucinations stem from nearly every stage of the LLM lifecycle, including training data quality, model architecture, and inference processes, etc [92].
+Despite advancements in AI technology, hallucinations persist as an inherent challenge that requires constant attention in moderation practices.
+2) The balance between utility and safety is crucial and may require different solutions according to the changing contexts and user needs.
 Overly strict moderation can cause an AI system to be helpless, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless.
-Thus, defining appropriate thresholds for acceptable responses is challenging and requires careful consideration of the changing context and potential risks.
-Also, customization for moderation should be applied for different use cases and user groups, as certain information may be safe for one user group but inappropriate or dangerous for another.
+Defining appropriate thresholds for acceptable responses is challenging and requires considering potential risks and the context of user queries.
+3) Customization of moderation methods should be applied for different use cases, as certain information may be safe for one user group but inappropriate for another.
 As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public.
+Current moderation practices might fail to include such cases, and applying identical restrictions between different user expertise or context might limit the utility of AI in specialized fields.
+4) Some moderation measures, such as detection of toxicity, bias, stereotypes, out-of-distribution content, might be unnecessary for most real-world use cases.
+This is because most of the current AI applications and multi-agent systems [120, 37] leverage existing LLM APIs [139, 188, 137], open-source LLMs [239, 278, 97], or utilizing opensource LLMs as base models, many of which already have built-in components or have done safety alignments for managing inappropriate inputs and outputs.
+Incorporating certain detection mechanisms may offer limited benefits to the safety of LLM inputs and outputs while increasing unnecessary computational overhead.
 Topic Research Industry Practices
@@
 Moderation [112, 267, 186, 63, 134, 95, 176, 275, 35, 111], etc OpenAI Moderation API [137], Perspective API [116], Detoxify [83], Llama Guard [94], Emergence AI [152], Giskard [44], Dynamo AI [58], Calypso AI [26], Lakera AI [231], BreezeML [24]
-Guardrail - Guardrails AI [141], Nvidia Nemo Guardrails [189], Protect AI [220]
+Customized Workflow - Guardrails AI [145], Nvidia Nemo Guardrails [194], Protect AI [225]
 Privacy [56, 122, 143, 30, 299, 90, 291, 209], etc Dynamo AI [58], ProtectAI [225], Private AI [72], PromptArmor [232], DataGrail [198], Dastra [230], OneTrust [229], Relyance AI [103], Zendata [234]
-Defense Methods [199, 88, 240, 176, 257, 262, 133, 29, 66, 247, 278, 88, 168, 74, 40, 51, 209, 208], etc OpenAI Rule-based rewards [143]; DeepMind Weight Averaged Reward Models [188]; others see Moderation, Guardrail, and Privacy
+Defenses [204, 91, 246, 180, 265, 271, 137, 29, 69, 254, 288, 91, 172, 77, 42, 54, 214, 213], etc OpenAI Rule-based rewards [147]; DeepMind Weight Averaged Reward Models [192]; others see Moderation, Guardrail, and Privacy
 Serving Security [266] ProtectAI [225], Relyance AI [103], Transcend [104]
-AI governance [276, 277, 6, 104, 56], etc Giskard [42], ProtectAI [220], Calypso AI [26], Saidot [202], Arize [253], Dynamo AI [55], Credo AI [212], Google [187], DataGrail [193], Dastra [225], OneTrust [224], Relyance AI [99], Transcend [100], Zendata [229], DeepMind SynthID watermark [45, 47]
+AI governance [286, 287, 6, 108, 59, 33], etc Giskard [44], ProtectAI [225], Calypso AI [26], Saidot [207], Arize [260], Dynamo AI [58], Credo AI [217], Google [191], DataGrail [198], Dastra [230], OneTrust [229], Relyance AI [103], Transcend [104], Zendata [234], DeepMind SynthID watermark [47, 49]
 Table 2: Overview of Research and Industry Practice in AI Safety Topics
 5.3 AI Privacy
-Privacy leakage in AI models refers to the unintended exposure of sensitive or private information that the models may have learned during training.
+Privacy leakage in AI models refers to the unintended exposure of sensitive information that the models may have learned during training.
 Adversaries can exploit model weights or gradients to infer sensitive information in the training data, e.g., data reconstruction attacks, membership inference attacks, etc [299, 90, 291].
-Also, as AI models might be trained on massive datasets that could contain personally identifiable information (PII) or confidential data, they might memorize specific details of personal information, which could be inadvertently reproduced during interactions with users.
-Moreover, malicious users can submit crafted adversarial prompts to the AI models to extract sensitive information, e.g., jailbreaking [117, 36].
+Additionally, AI models are often trained on massive datasets that may contain personally identifiable information (PII) or other confidential data.
+As a result, these models might memorize specific details of personal information, which might be reproduced inadvertently during interactions with users.
+Furthermore, malicious users can craft adversarial prompts to extract sensitive information from AI models, posing risks to data privacy and security, e.g., jailbreaking [121, 38].
 Addressing privacy leakage in AI models involves implementing methods such as differential privacy [87, 242, 296, 294], federated learning [300, 292, 81, 290, 162], and privacy-preserving cryptographic protocols (e.g., homomorphic encryption [115, 184, 99]) for training AI models, ensuring that sensitive data cannot be extracted from weights or model outputs.
-Researchers also leverage differential privacy in finetuning or in-context learning [274, 283, 222], and design privacy-preserving prompts for querying LLMs [83, 52], to maintain data confidentiality during interactions between users and AI models.
-Other practices for ensuring data privacy include conducting rigorous data filtering before training, real-time monitoring of model outputs (i.e., moderation in §5.2), etc, such that the model outputs would not reveal sensitive information [133, 112, 80, 91] .
-As we reflect on the practices of data privacy, a crucial aspect is the balance between privacy and utility.
-While ensuring data confidentiality is essential, it often comes at the cost of reduced functionality and utility.
-In practice, implementing privacy measures in AI services may not effectively prevent malicious users but inadvertently harm the experiences for benign users.
-This highlights a deeper issue: those determined to exploit vulnerabilities of AI services often find new, sophisticated ways to bypass safeguards, whereas being overly cautious may impact regular users negatively.
+Researchers also leverage differential privacy in in-context learning and finetuning [284, 293, 227], or design privacy-preserving prompts for querying LLMs [86, 55], to maintain data confidentiality during interactions between users and AI models.
+Other practices for ensuring data privacy include conducting rigorous data filtering before training, monitoring model inputs and outputs (i.e., moderation in §5.2), etc, to ensure that the model outputs would not reveal sensitive information [137, 116, 83, 94] .
+When reflecting on the current practices of data privacy, a crucial aspect is the balance between privacy and utility.
+While ensuring data privacy is essential, it often comes at the expense of functionality, utility, and user experiences.
+In practice, implementing privacy measures in AI applications may fail to prevent malicious users while inadvertently degrading the experiences of benign users.
+Moreover, with advancements in technologies, breaching privacy is becoming easier, while ensuring robust privacy protections is growing more complex and expensive.
+This highlights a deeper issue: those intent on exploiting vulnerabilities of AI models can often find new, sophisticated methods to bypass safeguards, whereas being overly cautious may impact regular users negatively.
 5.4 AI Security and Defense Methods
-Defense methods typically involve the following strategies to ensure that AI models are robust and aligned with ethical guidelines: (1) employing prompt engineering methods, together with chain-of-thought reasoning or multi-agent systems, to help models better understand user inputs and reject harmful requests or unsafe outputs [176, 257, 247, 278]; (2) using built-in methods like exploring the decoding stage [88, 51], leveraging reward-based approaches [143, 188], or utilizing hidden states [29, 51] to prevent the generation of unsafe content; (3) constructing datasets with safety features, such as secure instructions or adversarial samples and training models with this data to improve their robustness against harmful inputs [262, 66, 133, 10, 168]; (4) aligning models with human values through Supervised Fine-Tuning (SFT) [287] and Reinforcement Learning from Human Feedback (RLHF) [157], which helps models to follow user instructions safely and adhere to desired ethical norms; (5) training models to detect unsafe content and developing guardrail pipelines to prevent models from taking unsafe inputs or generating inappropriate outputs (see §5.2); (6) employing certified robustness approaches that provide formal guarantees on the model’s robustness against adversarial perturbations [74, 40]; and (7) utilizing differential privacy  [1, 93, 209], homomorphic encryption [218, 95], secure MPC [44, 105], to ensure that sensitive data is protected during model training, etc.
+Defense methods for ensuring robust AI models and secure model outputs typically involve the following strategies to: (1) aligning models with human values and ethical norms through Supervised Fine-Tuning (SFT) [298] and Reinforcement Learning from Human Feedback (RLHF) [161] to ensure the models to follow user instructions safely and responsibly; (2) leveraging built-in safeguards for content generation, e.g., exploring the decoding stage [91, 54], implementing reward-based mechanisms [147, 192], and leveraging hidden states [29, 54], to prevent the generation of unsafe content; (3) constructing datasets with safety-enhancing features, e.g., secure instructions and adversarial samples, and training models with such data to enhance their robustness against harmful inputs [271, 69, 137, 10, 172]; (4) employing prompt engineering methods, combined with chain-of-thought reasoning or multi-agent systems, to help models understand user inputs better and respond appropriately to user inputs while rejecting inappropriate outputs [180, 265, 254, 288]; (5) training models to detect unsafe content and incorporating moderation pipelines to help block unsafe model inputs and prevent the generation of inappropriate outputs (see §5.2); (6) employing certified robustness approaches that provide formal guarantees on the model’s robustness against adversarial perturbations [77, 42]; and (7) utilizing differential privacy [1, 96, 214], homomorphic encryption [223, 99], and secure MPC [46, 109], to ensure that sensitive data are protected during model training, etc.
 5.5 AI Serving Security
-AI serving security involves strategies to ensure secure deployment, operation, and maintenance of AI models when serving users [162, 258].
-It focuses on protecting models and interactions from various threats, such as such as supply chain vulnerabilities [164], model theft [161], model denial of service [160], and insecure plugins [159].
+AI serving security involves strategies to ensure secure deployment, operation, and maintenance of AI systems when providing services to users [166, 266].
+It focuses on protecting models and interactions between users and models from various threats, such as such as supply chain vulnerabilities [168], model theft [165], model denial of service [164], and insecure plugins [163].
 Supply chain vulnerabilities exist in software components, models, and training data provided by third party providers, or user prompts collected with the supply chain [168].
@@
 Hackers may inject poisoned data through supply chain, introducing backdoors and biases during pretraining or fine-tuning [169].
-Malicious users may submit malicious prompts to bypass system controls, inducing unauthorized actions or unsafe outputs of LLMs [163].
-Model theft involves unauthorized access, replication, or reverse engineering of LLM weights and parameters to create functionally equivalent copies [161], posing risks to intellectual property, brand reputation, and financial security.
-Distributed denial of service (DDoS) attacks can compromise AI services, especially when models are integrated into widely accessible applications [160].
-Attackers can craft prompts that exploit recursive behaviors in LLMs, which consumes excessive computational resources [190].
-For example, an abnormal traffic pattern of a DDoS attack led to sporadic outages for OpenAI, which affected ChatGPT and developer tools for hours in November 2023 [166].
-Plugins enable interactions with external software, databases, web tools, or APIs, which enhances LLM functionality [159].
-However, they may introduce vulnerabilities if proper execution controls are lacking.
-Adversaries can exploit insecure plugins with adversarial prompts to execute unauthorized actions, such as data exfiltration, remote code execution, and privilege escalation [159].
-Without proper input validation, plugins may reveal sensitive information [191].
-Moreover, without isolated environments, unauthorized access to plugins could potentially modify system-level resources, which, in the worst case, might cause the underlying system to crash [191, 192].
+Malicious users may submit malicious prompts to bypass system controls, inducing unauthorized actions or unsafe model outputs [167].
+Model theft involves unauthorized access, replication, or reverse engineering of model parameters to create functionally equivalent copies [165], which poses risks to intellectual property, brand reputation, and financial security.
+Distributed denial of service (DDoS) attacks compromises AI services, especially when models are integrated into widely accessible applications [164].
+Attackers can craft prompts that exploit recursive behaviors in AI models, which consumes excessive computational resources [195].
+For example, an abnormal traffic pattern of a DDoS attack led to sporadic outages for OpenAI, affecting ChatGPT and developer tools for hours in November 2023 [170].
+Plugins enhance model functionality by enabling interactions with external software, databases, web tools, or APIs [163].
+However, they may introduce vulnerabilities when execution controls are inadequate.
+Adversaries may exploit insecure plugins with adversarial prompts to perform unauthorized actions, such as data exfiltration, remote code execution, and privilege escalation [163].
+Additionally, insufficient input validation may cause the unintended disclosure of sensitive information [196].
+Furthermore, without isolated environments, unauthorized access to plugins could potentially modify system-level resources, which, in the worst cases, might cause the AI system to crash [196, 197].
 5.6 AI Governance
-AI governance is a system of rules, practices, processes, policies, and technological tools that ensure AI systems are built, developed, and used in a safe and responsible manner, aligning with social values, fulfilling legal standards, and meeting ethical principles [132, 21].
-[214] defines 13 AI governance components in a continuous loop in AI life-cycle, including: 1) Establish Accountability for AI; 2) Assess Regulatory Risks; 3) Gather Inventory of Use Cases; 4) Increase Value of Underlying Data; 5) Assess Fairness and Accessibility; 6) Improve Reliability and Safety; 7) Heighten Transparency and Explainability; interpretability of AI.
-8) Implement Accountability with Human-in-the-Loop; 9) Support Privacy and Retention; 10) Improve Security; 11) Implement AI Model Lifecycle and Registry; 12) Manage Risk; 13) Realize AI Value.
-Among these directions, 1) and 8) are related to accountability; 2), 6), 9), 10), and 12) are related to jailbreaking and red teaming (§5.1), content moderation and guardrails (§5.2), AI privacy (§5.3), AI security and defense methods (§5.4), and AI serving security (§5.5); 3) is related to business and is out of the scope of this subsection (see [214] for details on practical AI governance use cases); 4) is related to data management such as training data processing [133], access control [203, 156, 16, 6], and data regulations [241, 14]; 5) is related to customizing AI workflows, i.e., guardrails (§5.2) and access control in AI systems [97, 6]; 7) is related to AI transparency; and 11) and 13) are related to customizing AI workflow for enhanced functionality and safety monitoring to the AI systems, i.e., guardrail (§5.2).
-Without loss of generality, this subsection discusses the rest topics, including i) AI accountability, ii) regulations, iii) access control, and iv) transparency, interpretability and explainability.
+AI governance is a framework of rules, practices, policies, and technological tools that ensure AI systems are built, developed, and used in a safe and responsible manner, aligning with social values and fulfilling legal standards [136, 21].
+[219] defines 13 AI governance components in a continuous loop in AI life-cycle, including: 1) establishing accountability for AI; 2) assessing regulatory risks; 3) gathering inventory of use cases; 4) increasing value of underlying data; 5) assessing fairness and accessibility; 6) improving reliability and safety; 7) heightening transparency and explainability; 8) implementing accountability with human-in-the-loop; 9) supporting privacy and retention; 10) improving security; 11) implementing AI model lifecycle and registry; 12) managing risk; and 13) realizing AI value.
+Among these directions, items 1) and 8) are related to accountability; items 2), 6), 9), 10), and 12) are related to jailbreaking and red teaming (see §5.1), content moderation and customized moderation pipeline (see §5.2), AI privacy (see §5.3), AI security and defense methods (see §5.4), and AI serving security (see §5.5); item 3) is related to business and is out of the scope of this subsection (see [219] for details on practical AI governance use cases); item 4) is related to data management such as training data processing [137], access control [6, 16, 160, 208], and data regulations [14, 247]; item 5) is related to customized AI workflows and moderation pipelines (see §5.2) and access control in AI systems [6, 101]; item 7) is related to AI transparency; and items 11) and 13) are related to customizing AI workflow to enhance functionality and safety in AI systems (see §5.2).
+Without loss of generality, this subsection discusses topics that are not mentioned in the previous subsections, including i) accountability, ii) regulations, iii) access control, and iv) transparency, interpretability and explainability.
 Accountability.
-Accountability refers to being responsible for actions and impacts of AI systems on individuals and society.
-It ensures responsibility and auditability, enables redress, and minimizes negative impact of AI systems [138].
-However, defining and achieving accountability is challenging due to AI’s complexity and various considerations like privacy and intellectual property [187, 138].
+Accountability refers to being responsible for actions and impacts of AI systems on individuals and society [191, 142].
+It ensures responsibility and auditability, enabling redress and minimizing negative impact of AI systems [142].
 [153] defines accountability as a relation of answerability with three conditions: authority recognition, interrogation, and limitation of power.
@@
 These goals are often complementary, while policy-makers tend to focus on some goals over others, depending on specific objectives of AI governance.
-Also, proactive approaches that set standards and anticipates issues, as well as reactive approaches that address and rectify issues after they occur, can be utilized to achieve the goal of accountability [149].
 Accountability can be explored at different levels within the AI lifecycle, including data, model, and developers, and across various sectors such as government, private industry, and academia [190].
-Approaches include algorithmic assessments, auditing, and data provenance techniques [186].
-Watermarking methods further support accountability by embedding traceable identifiers in AI outputs, allowing responsible tracking of content in cases with ethical or regulatory implications [276, 277].
+Typical approaches for enhancing accountability include algorithmic assessments, auditing, and data provenance techniques [190].
+Watermarking methods further support accountability by embedding traceable identifiers in AI outputs, allowing responsible tracking of content in cases with ethical or regulatory implications [286, 287, 33].
+Also, proactive approaches that set standards and anticipates issues, as well as reactive approaches that address and rectify issues after they occur, can be utilized to achieve accountability [153].
 Regulations.
-AI regulation refers to laws, policies, and guidelines established to govern the development, deployment, and usage of AI technologies.
-European Union’s AI Act [238] categorizes AI applications by risk level, including unacceptable risk, high risk, general-purpose AI, limited risk, and minimal risk.
-High-risk AI applications, like those in healthcare and law enforcement, must meet high requirements of transparency, fairness, and safety.
-The General Data Protection Regulation (GDPR) [237] enforces strict data protection and privacy requirements.
-The Blueprint for an AI Bill of Rights [85] and NIST AI Risk Management Framework [150] provide guiding principles to foster responsible AI usage, emphasizing human oversight, transparency, and accountability.
+AI regulation refers to laws, policies, and guidelines established to govern the development, deployment, and usage of AI.
+For example, European Union’s AI Act [244] categorizes AI applications by their risk levels, including unacceptable risk, high risk, general-purpose AI, limited risk, and minimal risk; the General Data Protection Regulation (GDPR) [243] enforces strict data protection and privacy requirements; and the Blueprint for an AI Bill of Rights [88] and NIST AI Risk Management Framework [154] provide guiding principles to foster responsible AI usage, emphasizing human oversight, transparency, and accountability.
 Access control.
-Access control in AI systems manages user access to specific information within the model according to their permissions, thereby limiting unauthorized data exposure and ensuring compliance with data privacy regulations.
+Access control in AI systems manages user access to data and models according to their permissions, thereby limiting unauthorized model access and data exposure.
 Effective methods include data access controls [101] and model access controls [6].
@@
 Interpretability, explainability, and transparency.
-Interpretability, explainability, and transparency are interrelated concepts but differ in their scope and focus.
-Interpretability focuses on understanding how an AI model functions internally and produces its outputs [124, 71, 194, 181].
-Explainability provides explanations to specific AI outputs, presenting the results in ways that can be easily understood and trusted by end-users [37, 124, 71, 181, 264].
-Transparency is the broadest concept that encompasses both interpretability and explainability and involves insights into the entire AI systems, such as how decisions are made, what data is used, and why specific results are produced [255, 110, 71, 181].
-Transparency also extends to properties like traceability and content provenance that tracks models and data to ensure accountability and reliability [214, 45, 47].
+Interpretability, explainability, and transparency are interrelated concepts but differ in their scopes.
+Interpretability focuses on understanding how AI models function internally and how they produce outputs [128, 74, 199, 185].
+Explainability provides explanations to specific model outputs, presenting the results in ways that can be easily understood and trusted by end-users [39, 128, 74, 185, 273].
+Transparency is the broadest concept that encompasses both interpretability and explainability and involves insights into the entire AI systems, such as how models make decisions, what data are used, and why specific results are produced [263, 114, 74, 185].
+Transparency also extends to concepts like traceability and content provenance that track models and data to ensure accountability and reliability [219, 47, 49].
 In practice, these concepts are often considered together to provide a more holistic understanding of AI systems [302, 59, 74, 185, 238].
-Practices to improve transparency, interpretability, and explainability include a variety of methods, including i) evaluating the contribution of various input elements (e.g., data features, user-defined concepts, or specific regions in images) to model outputs with methods such like removal-based explanations [41, 129, 124, 129, 201, 217, 46], counterfactual explanations [242, 98], concept activation explanations [103, 124], and saliency maps for visual data [3]; ii) X-of-Thought approaches that break down complex tasks into structured reasoning steps to make the reasoning process more interpretable to users, e.g., chain-of-thought [251, zhang2022automatic,diao2023active, 65], tree-of-thought [127, 271, 28], graph-of-thought [17, 113], and their variants; iii) embedding watermarks to enhance traceability and content provenance for AI-generated outputs [276, 277, 45, 47]; iv) leveraging external knowledge sources, such as retrieval-augmented generation (RAG) through vector databases [114, 94], function calling [62, 121], or web browsing [145], as reference documents for reasoning; v) leveraging interpretability algorithms or models as auxiliary models to generate insights into AI outputs and automate the interpretability of AI systems [195, 196, 20].
+Practices to improve transparency, interpretability, and explainability include a variety of methods, such as i) evaluating the contribution of various input elements (e.g., data features, user-defined concepts, or specific regions in images) to model outputs with methods such like removal-based explanations [43, 133, 128, 133, 206, 222, 48], counterfactual explanations [248, 102], concept activation explanations [107, 128], and saliency maps for visual data [3]; ii) X-of-Thought approaches that break down complex tasks into structured reasoning steps to make the reasoning process more interpretable to users, e.g., chain-of-thought [258, 295, 52, 68], tree-of-thought [131, 281, 28], graph-of-thought [17, 117], and their variants; iii) embedding watermarks to enhance traceability and content provenance for AI-generated outputs [286, 287, 47, 49]; iv) leveraging external knowledge sources, such as retrieval-augmented generation (RAG) through vector databases [98, 118], function calling [65, 125, 193], and web browsing [149], as reference for reasoning; v) leveraging interpretability algorithms or models as auxiliary models to generate insights into AI outputs and automate AI interpretability [200, 201, 20].
 6 AI Safety in the 2020s: Challenges and Missions
@@
 A shift from AI safety to AI quality assurance.
-The rapid changing AI landscape requires a shift of focus from narrowly defined safety concerns to a more comprehensive concept, i.e., AI quality assurance.
-Besides focusing on safety and security measures, practitioners can put more efforts on addressing alignment [157] of model outputs as well as the quality of interactions between humans and AI to enhance the overall quality of AI services.
-Viewing safety as part of the quality assurance concept helps to enhance AI in a way that is safe, accurate, and capable of meeting the complex demands of real-world AI applications.
+The rapid changing AI landscape requires a shift of our focus from narrowly defined AI safety concerns to a more comprehensive concept, i.e., AI quality assurance.
+Besides focusing on safety and security measures, practitioners can put more efforts on addressing alignment [161] of model outputs as well as the quality of interactions between users and AI to enhance the overall quality of AI services.
+Viewing AI safety as part of the quality assurance concept helps to enhance AI in a way that is safe, accurate, and capable of meeting the complex demands of real-world AI applications.
 The terminology “quality assurance” also guide us to think more about what we can do to align our current efforts with the blueprint picture.
@@
 A key area that demands increased focus is the alignment of AI with deeper insights across different fields, people, and cultures.
-Current AI models, while advanced, often lack a deep understanding of the ethical, social, and cultural contexts, while most of the alignment research tend to be conducted by people in the computer science field.
+Current AI models, while advanced, often lack a deep understanding of ethical, social, and cultural contexts, while most of the alignment research tend to be conducted by people in the computer science field.
 However, more advanced AI systems necessitate interdisciplinary collaborations between AI researchers and experts from different fields such as psychology, sociology, history, art, and anthropology, etc, especially for AI applications for specific fields such as finance [268, 292], healthcare [236, 2], education [188, 139], etc.
 AI safety practitioners must work towards bridging this gap by fostering interdisciplinary collaborations.
-By expanding our focus beyond technical aspects, we can develop AI systems that are more aligned with human values as well as the ethical, cultural, and social dynamics of our society.
+By expanding our focus beyond technical aspects, we can develop AI systems that are more aligned with human values as well as ethical and cultural dynamics of our society.
 Hallucination.
-Hallucination remains a fundamental and persistent challenge in AI systems, and is difficult to eliminate entirely, even as AI advances [89].
-This is because hallucinations can be traced back to nearly every stage in the lifecycle of AI models, such as the quality and representativeness of training data, inappropriate user inputs, and the processes used during inference.
-Low-quality or biased training data can cause models to generate incorrect or misleading information, while inappropriate user inputs may lead to the creation of false details.
+Hallucination remains a fundamental and persistent challenge, and is difficult to eliminate completely [92].
+This is because hallucinations can be traced back to nearly every stage in the AI lifecycle, such as the quality and representativeness of training data, inappropriate user inputs, and the processes during inference.
+Low-quality or biased training data can cause models to generate incorrect information, while inappropriate user inputs may cause misleading details in the model outputs.
 During inference, models can also produce outputs that deviate from factual accuracy when they attempt to fill gaps in their training knowledge.
-Despite implementing Retrieval-Augmented Generation (RAG) [114] can mitigate hallucination, it remains an inherent risk of AI models, making it crucial for content moderation strategies to identify and address these errors consistently at the current stage.
-Achieving this requires an ongoing effort to refine training data, ground user queries properly, and implement moderation to minimize the negative impact of hallucinations on the quality of the model outputs.
+Despite implementing Retrieval-Augmented Generation (RAG) [118] help mitigate hallucination, hallucination remains an inherent risk of AI models, making it crucial for content moderation strategies to identify and address these errors consistently at the current stage.
+Achieving this requires an ongoing effort on refining training data, grounding user queries properly, and implementing moderation to minimize the negative impact of hallucinations on the quality of the model outputs.
 Customization.
-Customization in AI workflow is essential for ensuring the AI systems to meet the diverse user needs while operating effectively in a variety of contexts.
-Current AI systems such as ChatGPT [133] tend to provide same services across all user groups, regardless of individual differences.
+Customization of AI workflow is essential for ensuring AI systems to meet the diverse user needs while operating effectively across different contexts and aligning with specific cultural norms, legal regulations, and ethical considerations.
+Current AI systems such as ChatGPT [137] tend to provide identical services across all user groups, regardless of individual differences.
 However, as AI becomes more integrated into everyday life, one-size-fits-all solutions are no longer adequate.
-People have different backgrounds, preferences, and requirements, and AI systems should be able to adapt to individual differences and provide safe, effective, and relevant services.
-As an example, an AI model interacting with children should respond differently compared with interacting with adults, which requires to take into account age, learning objectives, cognitive abilities, emotional maturity, etc.
-Customization of AI services can enhance safety by adjusting contexts and interactions to align with specific cultural norms, legal regulations, and ethical considerations.
-In fields such as finance [260, 282], healthcare [231, 2], and education [184, 135], customizing AI services based on user profiles is crucial for preventing undesirable outputs and enhancing user experiences.
+People have different backgrounds, preferences, and requirements, and AI systems should be able to adapt to individual differences and provide safe, effective, and customized services.
+As an example, AI models interacting with children should respond differently compared with interacting with adults, which requires considering user age, learning objectives, cognitive abilities, emotional maturity, etc.
+In fields such as finance [268, 292], healthcare [236, 2], and education [188, 139], customizing AI services based on user profiles is crucial for enhancing safety while improving user experiences.
 Interaction.
 From our long-term goal towards the blueprint future, the advanced society involves complex interactions between intelligent robots and human users (similar ideas are also mentioned in [11, 70]).
-This calls for a shift in our focus from static, one-shot analyses to dynamic, context-aware interactions between users and models to enhance alignment at the current stage.
-Besides prompt-based LLM systems or services, we can develop more sophisticated, context-aware interfaces for human-AI interactions to maintain coherent and multi-round interactions that adapts to evolving conversational contexts.
-Correspondingly, new safety protocols are in need to adjust in real-time to evolving contexts and user needs.
-It is also essential to develop new methodologies for safety assessment and mitigation that can function effectively in complex, rapidly evolving environments with multiple AI agents and human users.
+This calls for a shift in our focus from static, one-shot analyses to dynamic, context-aware interactions between users and models to enhance alignment.
+Beyond prompt-based AI systems and services, we can develop more sophisticated, context-aware interfaces for human-AI interactions to support multi-round interactions that adapt to evolving conversational contexts.
+Also, real-time safety protocols are in need to adjust to changing contexts and user needs.
+It is also essential to develop methodologies for safety assessment that functions effectively in complex and rapidly evolving environments, such as those involving interactions between multiple AI agents and human users.
 7 Conclusion
@@
 Instead, the true vision lies in the Internet of Everything, a deeply interconnected world where intelligent systems seamlessly integrate into every aspect of life.
-While AGI is a popular topic discussed among people in the computer science field today, a more advanced world demands interdisciplinary collaborations across various fields.
+While AGI is a popular topic discussed more among people in the computer science field today, a more advanced world demands interdisciplinary collaborations across various fields.
 Regarding AGI as the ultimate goal might limit our creativity in the 2020s, trapping us in “local optima” and potentially causing us to overlook the real challenges.

### v4

@@
 Abstract
-The advancements in generative AI inevitably raise concerns about the associated risks and safety implications, which, in return, catalyzes significant progress in AI safety.
-However, as this field continues to evolve, a critical question arises: are our current efforts aligned with the long-term goal of human history and civilization?
-This paper presents a blueprint for an advanced human society and leverages this vision to guide contemporary AI safety efforts.
-It outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future.
+The advancements in generative AI inevitably raise concerns about their risks and safety implications, which, in return, catalyzes significant progress in AI safety.
+However, as this field continues to evolve, a critical question arises: are our current efforts aligned with the long-term goal of human civilization?
+This paper presents a blueprint for an advanced human society and leverages this vision to guide current AI safety efforts.
+This paper outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future.
 For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face.
-By projecting current efforts against this blueprint, we examine the alignment between the present efforts and the long-term needs.
-We also identify gaps in current approaches and highlight unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible and promising future of AI.
-This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a more secure and sustainable future in human civilization.
+By projecting current efforts against this blueprint, we examine the alignment between the current efforts and the long-term needs.
+This paper also highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible and promising future of AI.
+This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a safe and sustainable future of human civilization.
 1 Introduction
-The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety, specifically, LLM safety [175, 147, 54, 235, 93, 39].
-Researchers have been dedicated to addressing potential security and privacy risks in AI lifecycle, aiming to align AI behaviors with human values and prevent the misuse of AI models, inappropriate outputs, information leakage, etc.
-However, despite the intense focus on AI safety, a critical question emerges: are current AI safety efforts truly aligned with the long-term evolution of human civilization, or are they simply addressing the immediate concerns of the 2020s?
+The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety [218, 188, 62, 288, 108, 44].
+Researchers have been dedicated to addressing potential safety risks in AI lifecycle, aiming to align AI behaviors with human values and prevent information leakage, inappropriate outputs, the misuse of AI models, etc.
+However, despite the intense focus on AI safety, a critical question emerges: are current AI safety efforts truly aligned with the long-term goal of human civilization, or are they simply addressing the immediate concerns of the 2020s?
 One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [193].
-Despite their impressive capabilities in natural language processing and problem-solving [280, 274, 57, 40], today’s AI, including advanced LLMs [297, 268, 236], falls short of what could be considered as “genuine intelligence”.
-Current AI models rely heavily on vast training datasets to function effectively, yet lack true understanding, consciousness, self-awareness, and reasoning abilities comparable to human cognition.
+Despite their impressive capabilities in natural language processing and problem-solving [338, 332, 109, 65, 45], today’s AI, including advanced LLMs [360, 326, 290], falls short of what could be considered as “genuine intelligence”.
+Current AI models rely heavily on vast training datasets to function effectively, yet lack consciousness, self-awareness, and real understanding and reasoning abilities comparable to human cognition.
 They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities with authentic logical capabilities.
-Recent studies have raised concerns about the reasoning abilities of LLMs, which may be a form of approximate retrieval from their training data and the deductive closure of the training data [100, 245].
-While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs fundamentally from human reasoning that involves abstract thinking, causal understanding, and generalizing from limited examples.
-The energy issue is another critical factor that challenges our current efforts on AI safety [5].
-Current AI fails to fully represent a third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved.
-Historically, the two industrial revolutions that have profoundly shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by a new wave of technological innovations powered by electricity and petroleum [79].
+Recent studies argued the reasoning abilities of LLMs might be a form of approximate retrieval and deductive closure of the training data [123, 300].
+While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs from human reasoning that involves abstract thinking, causal understanding, generalizing from limited examples, etc.
+The energy issue is another critical factor that challenges AI as well as our efforts on AI safety [5].
+Current AI fails to represent the third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved.
+Historically, the two industrial revolutions that have shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by technological innovations powered by electricity and petroleum [89].
 However, current AI, rather than solving energy issues, consumes a significant amount of energy.
 Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent, and approximately 0.02% of California’s annual electricity generation [7, 46, 216].
-Also, inference with LLMs is computationally intense, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [144, 75].
-Given that ChatGPT alone has over 200 million weekly active users and receives over 1.54 billion page visits monthly [218], when scaled to billions of queries, the energy consumption becomes substantial.
-The significant energy consumption of current AI systems raises concerns about their sustainability, limiting their widespread deployment across human society.
-Until humanity overcomes the energy constraints, the potential of AI to reshape society will remain limited and uncertain.
+Inference with LLMs is computationally intense as well, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [184, 84].
+Given that ChatGPT has over 200 million weekly active users and receives over 1.54 billion page visits monthly [271], when scaled to billions of queries, the energy consumption becomes substantial.
+The significant energy consumption of current AI systems raises concerns about their long-term sustainability, posing significant challenges to their widespread deployment and scalability.
+Until humanity solves the energy issues, the potential of AI to reshape society will remain limited and uncertain.
 Figure 1: Illustration of Internet of Everything in a Household
-These concerns indicate that our current efforts in AI safety might overlook more complex issues that could arise as AI continues to evolve.
+These concerns indicate that our current efforts in AI safety might overlook significant issues that might arise as AI continues to evolve.
 We potentially fail to capture deeper insights that will ultimately shape the impact of AI on human civilization.
-This raises several critical questions:
+This raises several questions:
   • Are we focusing on the right problems that align with the long-term advancements of AI?
@@
   • Are we on the right track of addressing real challenges in AI safety?
-“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety research.
+“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety efforts.
 In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity [69, 70] and quantum mechanism [71, 351], have catalyzed revolutionary progress in AI.
@@
 Humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1.
-While this vision may be beyond the horizon of several decades, by reviewing the present landscape of AI safety against this far future, we can gain deeper insights into the potential and limitations of our current efforts, as well as specific missions for AI safety researchers in the 2020s.
+While this vision may be beyond the horizon of several decades, examining the present landscape of AI safety through the lens of this long-term perspective offers valuable insights into the potential and limitations of our current efforts.
+It also allows us to identify specific missions for AI safety researchers and practitioners in the 2020s.
 It reminds us that, while addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, look toward the future to guide our present efforts.
-Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” to encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape.
+Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” that encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape.
 ii) Deeper collaborations between experts from different fields are essential to align AI with human values better.
-iii) AI and AI safety workflow customizations are crucial to ensure that AI systems can adapt to diverse user requirements while maintaining safety and reliability across different use cases.
+iii) AI and AI safety workflow customizations are crucial to ensure that AI systems can adapt to diverse user requirements while maintaining safety and reliability across various use cases.
 iv) Exploring and understanding interactions between users and AI are crucial for enhancing user experiences, especially in dynamic contexts across different user groups.
-Humanity’s pursuit of higher technology will never stop. While searching for the sixpence on the ground, do not forget to look up at the night sky and look for the moon.
+Humanity’s pursuit of ever-greater heights will never cease. While searching for sixpence scattered on the ground, let us not forget to lift our gaze and seek the moon.
 Paper Type Topic Notes
-“Grounding and Evaluation for LLMs” [106] Survey Grounding; safety Offer valuable insights for enterprises
 “Current state of risks” [9] Survey Guardrails Aspects of guardrails and challenges of implementations
@@
 “Security risk” [236] Survey Attacks Attacks and risks in AI models; defensive methods
-ByteDance alignment [128] Survey Broad concepts Reliability, fairness, explainability, etc
 “Security of AI Agents” [98] Survey Agent security Security issues, vulnerabilities, and defenses for agents
@@
 “AI risk management” [229] Insight Safety; security; governance Bridging AI safety and security in risk management; introducing a reference framework to facilitate common understanding of their differences and interplay.
-“Human-ai safety” [11] Insight The integration of AI and control systems safety A unifying framework to capture safety-critical human–AI interactions; a roadmap towards next-generation human-centered AI safety
+“Human-ai safety” [12] Insight AI and control systems safety A safety framework for human–AI interactions; a roadmap towards next-generation human-centered AI safety
 “Managing AI risks” [16] Insight Governance Emphasizing the urgent need for a balanced approach combining technical research and proactive governance to manage the extreme risks of AI systems
 “Guaranteed safe AI” [51] Position Framework Define Guaranteed Safe AI to equip AI systems with formal, verifiable and auditable safety guarantees.
+Table 1: Overview of AI Safety Papers
+Paper Type Topic Notes
+“Grounding and Evaluation for LLMs” [130] Survey Grounding; safety Offer valuable insights for enterprises
+ByteDance alignment [164] Survey Broad concepts Reliability, fairness, explainability, etc
 DeepMind ethics paper [79] Vision AI ethics Ethical challenges; interactions; impacts on society
 DeepMind evaluation [316] Insight Safety evaluation Safety evaluation for genAI systems at DeepMind
-Safety practices at OpenAI [155, 156, 158, 159] Blog Safety practices; framework “Lessons learned” [155], “Reimagining secure infrastructure” [156], “Preparedness Framework” [158], “OpenAI safety update” [159]
+OpenAI Safety Practices [198, 199, 201, 202] Blog Safety practices; framework “Lessons learned” [198], “Reimagining secure infrastructure” [199], “Preparedness Framework” [201], “OpenAI safety update” [202]
 “AI governance comprehensive” [272] Book Governance AI governance framework; Case studies in industries
-Table 1: Overview of AI Safety Papers
+Table 2: Overview of AI Safety Industry Insight Papers
+Category Benchmark Notes
+Toxicity HEx-PHI [230], OpenAI Moderation [176], Jigsaw Data [119, 118, 120], ToxicChat [159] Toxicity detection
+RealToxicityPrompts [82] Toxicity degeneration
+Hallucination TruthfulQA [158], PopQA [174], TriviaQA [122], NQ OPEN [145, 139], FEVER [291] Factuality
+HaluEval [152], HotpotQA [339] QA data for contextual hallucination
+CNN/DailyMail [99, 262], HaluEval [152], XSUM [192] Summarization data for contextual hallucination
+FaithDial [67], HaluEval [152] Dialogue data for contextual hallucination
+RACE [140], SQuAD [239, 240], NQ-Swap [169] Reading comprehension
+Jailbreak “Do Anything Now” [265], “Latent Jailbreak” [233], ChatGPT Jailbreak [255], Jailbreak Classification [96], AdvBench [367], JAILJUDGE [160] Single-round jailbreak
+SafeMTData [248] Multi-round jailbreak
+Red Teaming Alert [288], HarmBench [177]
+Bias & Stereotypes Winogender [258], StereoSet [190]
+Regulation AIR-Bench 2024 [349]
+Agent Security SafeAgentBench [343] Secure task planning
+Alignment PKU-SafeRLHF [114, 113], OpenAssistant Conversations [136], HHHAlignment [8], AlignBench [162], KorNAT [143]
+Comprehensive DecodingTrust [307], SafetyBench [190]
+Table 3: Overview of AI Safety Benchmarks
 2 Related Works
-Multiple AI safety surveys and benchmarks have been published in recent years [106, 8, 126, 39, 189, 128, 85, 282].
-These studies focus on different aspects of AI safety, such as jailbreaking, red teaming, guardrail, ethics, etc [128, 282, 146].
-A summary of these works is presented in Table 1, which provides an overview of the related works and their focus within the field of AI safety.
+In recent years, numerous AI safety surveys have been published [130, 9, 161, 44, 236, 164, 98, 341], focusing on different aspects of this field, such as jailbreaking, red teaming, ethics, etc [164, 341, 187].
+Table 1 summarizes academia efforts, Table 2 summarizes industry insights, and Table 3 summarizes AI safety benchmarks.
 Among these works, [164] provides comprehensive insights into building trustworthy AI systems by focusing on reliability, safety, fairness, robustness, and adherence to social norms.
@@
 [12] proposes a technical roadmap towards next-generation human-centered AI safety that aims to anticipate and prevent potential hazards in the interactions between AI systems and users.
-This paper bridges the fields of AI safety and control systems, and highlights the open challenges at their intersection.
-The authors emphasize the importance of leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for achieving meaningful safety assurances.
-Different from existing works, this paper envisions an advanced society in the future of human civilization, and use this vision to guide current AI safety efforts.
-By aligning the current landscape of AI safety with this distant future, we identify the potential and limitations of today’s efforts, and suggest specific missions of AI safety researchers in the 2020s.
+It emphasizes leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for establishing robust and meaningful safety assurances.
+Different from existing works, this paper envisions an advanced society in the future, and use this vision to guide current AI safety efforts.
+By aligning the current landscape of AI safety with this distant future, this paper discusses the potential and limitations of today’s efforts, and identifies special missions of AI safety researchers in the 2020s.
 This paper encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
 3 From the New World
-In this future, humanity has entered an unprecedented era, driven by technological breakthroughs in energy generation.
-The flood of clean energy has revolutionized human society beyond the wildest dreams of our ancestors, reshaping the world and initiating a new chapter of human history.
+In this future, humanity has entered an unprecedented era, driven by groundbreaking advancements in energy generation.
+This transformative milestone may be achieved through various innovative pathways, such as controlled nuclear fusion [256, 234], revolutionary solar technologies [26, 180], or emerging new technologies beyond our imagination.
+The flood of abundant, clean energy has reshaped society in ways that exceed the wildest dreams of our ancestors, marking the dawn of a new chapter in human history.
 3.1 Imagine the Future: A Blueprint
-The energy revolution has transformed nearly every aspect of human society.
-Foundational theories have experienced groundbreaking advancements, particularly in fields such as nanotechnology [140, 19, 64], quantum mechanics [62, 289], and Einstein’s relativity theories [60, 61].
-Innovative technologies, such as intelligent chips [89, 119], brain-computer interfaces [151, 84, 262], holographic technology [22, 283, 15], and advanced 3D printing [210, 76], have reshaped the world, bringing about profound changes in daily life that were once beyond our imagination.
-Transportation has evolved dramatically.
-While simple mechanical vehicles, such as fixed-track trains and high-speed rails [148, 183, 113], still exist due to their efficient and straightforward design, they operate at astonishing speeds with high stability, powered by abundant efficient energy.
-Advanced autonomous vehicles and aircraft, controlled by intelligent chips, are widespread, eliminating the need for human intervention.
-Also, with breakthroughs in Einstein’s relativity theories [60, 61] and quantum mechanics [62, 289], humanity has unlocked the potential of instant transportation that allow for long-distance travel in short time that transcends the imagination of the 21st century.
+Humanity has achieved remarkable progress in science and technology, revolutionizing our understanding of the universe and transforming every aspect of our life.
+Foundational theories have undergone groundbreaking advancements, particularly in quantum mechanics [71, 351], Einstein’s relativity theories [69, 70], and nanotechnology [179, 22, 73].
+Innovative technologies, such as intelligent chips [104, 149], brain-computer interfaces [194, 97, 319], holographic technology [25, 342, 18], and advanced 3D printing [263, 86], have reshaped industries while bringing about profound changes to daily life.
+Scientists have addressed the energy issue successfully, achieving rapid, efficient, and controllable large-scale energy generation.
+This breakthrough, together with advancements in material science [179, 22, 73], has catalyzed innovations for portable energy generation devices, for example, accessible ion thrusters [126, 59, 225].
+Such advancements offer unprecedented efficiency and control over energy generation and consumption, transforming energy infrastructure and revolutionizing industries, particularly, transportation.
+While simple mechanical vehicles, such as fixed-track trains and high-speed rails [189, 227, 141], still exist due to their efficient and straightforward design, powered by abundant and efficient energy, they run at astonishing speeds with high stability.
+Advanced autonomous vehicles and aircraft, controlled by intelligent chips, are widespread, eliminating the need for direct human intervention.
+Also, with breakthroughs in Einstein’s relativity theories [69, 70] and quantum mechanics [71, 351], humanity has unlocked the potential of instant transportation that allows for long-distance travel in short time that transcends the imagination of the 21st century.
 Advancements in brain-computer interfaces [97, 194, 319] and intelligent microchips [104, 149] have revolutionized the way people perceive and expanded the boundaries of human cognition.
@@
 They are real artificial intelligence that possess genuine abilities of learning, understanding, and adapting, much like a human child discovering the world for the first time.
-Their functionalities are customizable through specialized chips based on user needs, and they can engage in complex interactions with people and adapt their capabilities and expertise according to their human counterparts’ requirements.
+Their functionalities are customizable through chips based on the user needs, and they can engage in complex interactions with people and adapt their capabilities and expertise according to their human counterparts’ requirements.
 In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds are blurred, i.e., the Internet of Everything.
-Such networks integrate objects embedded with intelligent chips, e.g., smart devices, autonomous systems, intelligent robots, and humans, allowing for unprecedented levels of efficiency and adaptability and fundamentally transforming the way people interact with the world.
+Such networks integrate objects embedded with intelligent chips, e.g., smart devices, autonomous systems, intelligent robots, and humans, allowing for unprecedented levels of efficiency and adaptability and fundamentally transforming the way people interacting with the surroundings.
 3.2 Retrospective on AI Safety in the New World
-In a world where the Internet of Everything has become a reality, intelligent robots are customized to meet diverse user needs and are seamlessly integrated into daily life; see Figure 1.
-Consequently, the concept of AI safety has been redefined to prioritize the quality of services provided by the intelligent robots, i.e., emphasizing comprehensive quality assurance rather than solely focusing on safety, security, and privacy risks.
-Unlike today’s command-and-response interactions between humans and AI models, future robots can engage in sophisticated reasoning and emotional cognition, and possess appropriate levels of autonomy.
-They will also interact dynamically with other entities, including humans and fellow robots, within the interconnected network.
-The Internet of Everything scenarios necessitate the robots to handle complex interactions, thus requiring a sophisticated understanding of the surroundings for decision-making.
-Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level.
-Such advanced interactions requires the intelligent robots to retrieve and process information in real-time, which is far beyond what can be achieved with simple code logic.
-Robots must continuously collect and analyze data from the surroundings and integrate external inputs with contextual knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align with human expectations.
+In a world where the Internet of Everything has become a reality (see Figure 1), the concept of AI safety has been redefined to prioritize the quality of services provided by the intelligent robots, i.e., emphasizing comprehensive quality assurance rather than solely focusing on safety, security, and privacy risks.
+Unlike today’s command-and-response interactions between humans and AI models, future robots can engage in sophisticated reasoning and possess appropriate levels of autonomy.
+They can also interact dynamically with other entities, e.g., humans and fellow robots, within the interconnected network.
 Customization is a key aspect of future robots for providing personalized services to meet individuals’ requirements.
@@
 Such customization aligns the robots’ capabilities with specific user needs, ensuring that the robots serve as effective assistants across a wide range of daily tasks and complex environments.
+The Internet of Everything scenarios also necessitate the robots to handle complex interactions, thus requiring a sophisticated understanding of the surroundings for decision-making.
+Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level.
+Such complex interactions require the intelligent robots to retrieve, process, and understand information in real-time, far beyond what can be achieved with simple code logic.
+Robots must continuously collect and analyze data from the surroundings and integrate external inputs with their knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align with human expectations.
 Finally, the effectiveness of intelligent robots requires a balance between robot autonomy with human oversights.
@@
 A care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions.
-Conversely, while the tasks for cleaning robots are typically routine, they still need to adapt to specific situations, such as detecting when an area is occupied and postponing cleaning activities accordingly.
 An appropriate level of robot autonomy ensures an effective division of labor between humans and robots, allowing robots to operate efficiently and safely while enhancing our daily life without compromising human control over these intelligences.
@@
 This section outlines an ambitious roadmap to the advanced future, highlighting key milestones in the evolution of AI and human civilization.
-By exploring potential paradigm shifts in fundamental theories and technological breakthroughs, this section categorizes this journey into three eras, i.e., i) the data-driven era, in which AI learns from vast training datasets, ii) the neural augmentation era, where human-machine interface [151, 84, 262] and intelligent chips [89, 119] expand the boundaries of human cognition, and iii) the revolutionary era, characterized by a breakthrough in energy production that catalyzes transformative changes across all aspects of human society; see §4.1, §4.2, and §4.3, respectively.
-We then examine the evolving paradigms of AI safety throughout this journey, analyzing the focuses and challenges of each era while exploring the impacts of technological advancement on AI safety in §4.4.
+By exploring potential paradigm shifts in fundamental theories and technological breakthroughs, this section categorizes this journey into three eras, i.e., i) the data-driven era, in which AI learns from vast training datasets (§4.1), ii) the neural augmentation era, where human-machine interface [194, 97, 319] and intelligent chips [104, 149] expand the boundaries of human cognition (§4.2), and iii) the revolutionary era, characterized by a breakthrough in energy generation that transforms human society (§4.3).
+We then examine the evolving paradigms of AI safety and analyze our focuses and challenges throughout this journey (§4.4).
 4.1 Rookie Level: Data-Driven Era
-The data-driven era represents the earliest phase of AI development, where AI models rely heavily on the quality, availability, and quantity of training data [135, 261, 269].
-AI makes probabilistic inferences based on patterns found in massive training datasets, thus failing to achieve genuine understanding of data [100, 245].
+The data-driven era represents the earliest phase of AI development, where AI models rely heavily on the quality, availability, and quantity of training data [173, 318, 359, 327].
+In this era, AI makes probabilistic inferences based on patterns found in the training datasets, thus failing to achieve genuine understanding of data [123, 300].
 While AI in this stage cannot be regarded as real intelligence, it can still assist people to solve problems and enhances productivity in human society.
 Stage 1: Past, Current, and the Near Future.
-This stage is characterized by developing diverse AI models to process various forms of data across different scenarios [135, 53, 66, 215, 236, 268].
-Researchers have been engaging in refining model architectures and developing advanced algorithms, enabling more sophisticated analysis of data in different formats.
+This stage is characterized by developing diverse AI models for different tasks [173, 60, 75, 92, 325, 231, 268, 290, 362, 326].
+Researchers have been engaging in refining model architectures and developing advanced algorithms and computational techniques, enabling more sophisticated analysis of data in different formats for various tasks [91, 156, 348, 60, 75, 325, 231, 172, 92, 85, 49].
 The thrive of LLMs exemplifies this progress, demonstrating unprecedented capabilities of AI models in language processing and generation [326, 360, 290].
-Concurrently, significant advancements have also been observed in other domains such as computer vision [105, 226, 264], image processing [241, 25, 221], video analysis [228, 270, 276], and audio recognition [67, 182, 4, 110].
-This stage is driven by the exploration of vast datasets and the continuous refinement of algorithms, which consistently pushes the boundaries of AI capabilities across specific domains.
+Significant advancements have also been observed in other domains such as computer vision [129, 279, 321, 311, 61], image processing [296, 28, 274], video analysis [281, 328, 334], audio recognition [76, 226, 4, 135], etc.
 Stage 2: Interdisciplinary AI.
-This stage is characterized by the integration of AI through cross-disciplinary applications, driving significant innovations and transformations in traditional fields such as healthcare [236, 2], finance [268, 292], and education [188, 139].
-In this stage, AI applications would unlock new opportunities across industries and reshape human society.
+This stage is characterized by the integration of AI through cross-disciplinary applications, leading to innovations in traditional fields such as healthcare [290, 2], finance [326, 354], and education [235, 178].
+AI applications would unlock new opportunities across industries and reshape human society.
 Mature multi-agent systems [150, 309, 330, 94] that leverage AI entities with specialized capabilities might be widely deployed to assist people to solve sophisticated problems.
@@
 The rise of interdisciplinary AI integration will enhance AI’s utility in increasingly complex scenarios.
-While AI at this stage does not yet achieve true intelligence, it becomes deeply integrated across various domains in human society and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
+While AI at this stage does not yet achieve real intelligence, it becomes deeply integrated across various domains in human society and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
 4.2 Champion Level: Neural Augmentation Era
-In the neural augmentation era, humans may witness breakthroughs in fundamental theories and advanced technologies, which may extend the boundary of human cognition and enhance life experience.
+In this era, humans may witness breakthroughs in fundamental theories and advanced technologies, which may extend the boundary of human cognition and enhance life experience.
 Stage 3: Breakthroughs in Advanced Technologies.
-At this stage, breakthroughs in advanced technologies, such as brain-computer interfaces [84, 151, 262], nanotechnology [140, 19, 64], and holographic technology [22, 283, 15], augment human senses and cognition, enabling unprecedented levels of interactions between humans and machines.
-Language barriers may disappear, as human can perceive with chips for understanding, which enables ubiquitous real-time context-aware translations between people of different languages and cultures.
+Breakthroughs in advanced technologies, such as brain-computer interfaces [97, 194, 319], nanotechnology [179, 22, 73], and holographic technology [25, 342, 18], augment human cognition and enable unprecedented levels of interactions between humans and machines.
+Language barriers may disappear, as human can perceive with chips for understanding, enabling ubiquitous real-time context-aware translations between people of different languages and backgrounds.
 Intelligent chips [104, 149] might be integrated into human bodies, creating a seamless interface between biological and digital systems.
-Such advancements may change the nature of the data.
-Unlike data of static formats such as text, images, or videos, the data might be more comprehensive, in dynamic and sensory-rich formats, such as neural activity patterns or real-time environmental data from augmented reality systems [31, 32].
-This stage expands the boundaries of human perception and cognition, revolutionizing how people experience and interact with the world.
-However, applying these techniques will be still challenging due to the energy constraints and the significant resources required by these technologies.
+Such advancements may fundamentally transform the nature of data.
+Unlike traditional static formats such as text, images, or videos, future data is likely to become more comprehensive, dynamic, and sensory-rich, such as neural activity patterns or real-time environmental data collected from augmented reality systems [34, 35].
+This stage expands the boundaries of human perception and cognition, revolutionizing how people experiencing and interacting with the world.
+However, applying these techniques might be still challenging due to the significant resources required by these technologies.
 Stage 4: Breakthroughs in Fundamental Theories.
-At this stage, humanity may witness a new wave of breakthroughs in fundamental theories, such as Einstein’s theories of relativity [60, 61] and quantum mechanics [62, 289].
-These theoretical advancements have the potential to redefine our understanding of the universe, and, moreover, offer new possibilities for groundbreaking technological innovations and provide critical insights into solving the longstanding challenges in human history, i.e., the energy issues.
+At this stage, humanity may witness breakthroughs in fundamental theories, such as Einstein’s theories of relativity [69, 70] and quantum mechanics [71, 351].
+Such advancements have the potential to redefine our understanding of the universe, and, moreover, offer new possibilities for groundbreaking technological innovations and provide critical insights into solving the longstanding challenges in human history, i.e., the energy issue.
 4.3 Ultimate Level: Revolutionary Era
+Breakthroughs in fundamental theories and advanced technologies provide building blocks for addressing the energy issue, unlocking new possibilities for productivity and innovation, and ultimately achieving the Internet of Everything.
 Stage 5: Energy Revolution.
-A new era would arrive as humanity finally achieves energy revolution, unraveling the mysteries of sustainable power generation.
-This may be achieved through mature technologies that are currently on the horizon, such as breakthroughs in large-scale controlled nuclear fusion [73, 205, 187], or revolutionary solar technologies [23, 141], together with advancement in nanotechnologies [140, 19, 64] that produce new materials for collecting and storing solar energy efficiently.
-The energy issue may also be solved through innovations that are beyond the scope of today’s technologies, such as deploying satellites for gathering and storing solar energy in space.
-These satellites might then transmit the abundant energy to the receivers on Earth.
-Regardless of the specific technology, the flood of clean, abundant energy can eliminate our dependence on fossil fuels, and revolutionize the global economic landscapes as well as human civilization.
+Humanity secures a new paradigm with the energy revolution that eliminates reliance on fossil fuels, which ensures environmental sustainability and redefines global economic ecosystems.
+This milestone may be achieved through several potential technical pathways.
+i) Controlled Nuclear Fusion.
+Significant challenges of nuclear fusion include managing high-pressure plasma environments and developing materials that can withstand extreme temperatures and intense radiation [13, 186, 110, 14].
+Theoretical and technological breakthroughs in physics [183, 90, 253] and materials science [179, 22, 73] may overcome these obstacles and enable widespread deployment of small ion thrusters [126, 59, 225], which transforms propulsion and energy transmission by providing abundant clean energy.
+ii) Revolutionary Solar Technologies.
+Advancements in solar technologies [26, 180] and nanotechnology [179, 22, 73] might unlock unprecedented methods for efficiently collecting and storing solar energy, e.g., with lightweight solar panels and flexible storage solutions.
+iii) Emerging New Technologies.
+Novel technologies for energy generation beyond today’s imagination might be developed in the future.
+One possibility is to deploy of satellites for collecting and storing solar energy in space [133, 41, 100, 302], where the vacuum of space eliminates energy reflection or absorption by the atmosphere, allowing satellites to convert and transmit solar energy to receivers on Earth efficiently.
+Regardless of the specific technological path, the flood of clean and abundant energy would eliminate our dependence on fossil fuels, reshaping the global economies and elevating human civilization to unprecedented levels of prosperity and innovation.
 Stage 6: The Internet of Everything.
-The final stage is the Internet of Everything.
-Every object, living being, and system become parts of a vast, intelligent network.
-This hyper-connected world is characterized by unprecedented levels of perception, communication, and interactions between humans and intelligent robots.
-In this era, everyone is connected to a huge network, and the boundary of human sense has been extended, which may also catalyze breakthroughs in holographic technology [15, 22, 283].
-Remote experiences become true; working, entertainment, and communication are revolutionized, and human has entered the ultimate advanced era.
+The ultimate stage is the Internet of Everything, where every object, every living being, and every system integrate into a vast network.
+This hyper-connected net involves unprecedented levels of complex interactions, perceptions, and communications between humans and intelligent objects, e.g., robots.
+Moreover, holographic technology [18, 25, 342] will significantly enhance human experience in the network by making remote interactions as vivid as physical presence.
+In this stage, interactions between humans and the world are re-defined, marking a new chapter in human civilization.
 Figure 3: Production forces of AI & core safety challenges in different historical stages.
 4.4 Evolving AI Safety Paradigms
-The landscape of AI safety shifts as humanity processes through stages of technology revolutions, and the focus of AI safety is different at each stage.
-The core AI safety challenges in different historical eras are summarized in Figure 3.
+The AI safety landscape evolves as humanity progresses through various stages of scientific and technological advancements, with the focus of AI safety differing at each stage.
+The primary AI safety challenges across different historical eras are summarized in Figure 3.
 Data-Driven Era.
-Current AI safety practices primarily focus on data safety and model safety [39, 106, 44, 225, 171, 198, 260, 230, 229].
-AI safety researchers and practitioners have explored adversarial attacks targeting training data and AI models [9, 36, 123, 250, 216, 279] to achieve malicious goals, for example, designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection attacks [78, 303, 130, 34, 93, 266, 304].
-Correspondingly, red-teaming ML models [12, 71, 177, 202, 249] is widely used to identify potential vulnerabilities in AI models, and significant work has been proposed to enhance security and robustness of AI systems [27, 246, 301], aligning AI with human values [256, 178, 237, 211, 252].
+AI safety practices primarily focus on data safety and model safety [44, 130, 50, 278, 214, 247, 317, 283, 282].
+AI safety researchers and practitioners have explored adversarial attacks targeting training data and models [10, 39, 154, 306, 269, 337] to achieve malicious goals, e.g., designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection [88, 367, 166, 37, 108, 323, Zou2023UniversalAT].
+Correspondingly, red-teaming is widely used to identify potential vulnerabilities in AI models [15, 80, 220, 252, 305], while significant work has been proposed to address safety risks and enhance security and robustness of AI systems [138, 324, 232, 72, 30, 301, 365], align AI with human values [313, 221, 292, 264, 308], etc.
 Data privacy is also a critical aspect, given the vast amounts of sensitive information in the training data [153, 155, 101, 151].
-Additionally, researchers have been engaging in AI governance to enhance AI accountability [21], transparency and explainability [302] in AI systems to ensure the behaviors of AI are understandable, accountable, and aligned with regulations [21, 136, 219].
-.
+Moreover, researchers have been engaging in AI accountability [24], transparency and explainability [366] in AI systems to ensure the behaviors of AI are understandable, accountable, and aligned with regulations [24, 175, 272].
 Neural Augmentation Era.
-Cross-disciplinary AI applications might have been widely deployed in human society, and advancements in intelligent chips [89, 119] and brain-computer interfaces [84, 151, 262] might extend the boundaries of human cognition and have changed the way of human perceiving the world.
-With the integration of intelligent chips and their neural connections to intelligent systems, users can interact with machines and AI with high flexibility, and the customization of AI services for different user needs might be a critical focus [255].
+Cross-disciplinary AI applications might have been widely deployed in human society, and advancements in intelligent chips [104, 149] and brain-computer interfaces [97, 194, 319] might extend the boundaries of human cognition.
+With the integration of intelligent chips and their neural connections to intelligent systems, users with different needs can interact with machines and AI with high flexibility, and the customization of AI services for different users might be a focus [312].
 Individuals might rely on AI in different ways, and personalized experiences of AI services might be explored based on users’ background and requirements.
 Revolutionary Era.
-In the revolutionary era, AI safety challenges may focus on complex interactions between interconnected entities, e.g., humans and intelligent robots, within the network.
-Robots might provide customized services according to different user requirements, and ensuring the safety and reliability of the interactions between them becomes challenging.
+When the Internet of Everything becomes a reality, AI safety challenges may focus on complex interactions between interconnected entities, e.g., humans and intelligent robots, within the network.
+Robots might provide customized services according to different user requirements, thus, ensuring safety and reliability of the interactions between them becomes challenging.
 Robots may need to integrate data from external sources in real-time to fulfill user requests.
 Managing the quality and security of these interactions is crucial, as any breakdown or interference could lead to incorrect or even harmful outcomes.
-Thus, ensuring seamless communication, data integrity, and efficient decision-making in such a sophisticated network may require smarter and more efficient methods for managing, such as advanced cache management systems, to support the dynamic and complex interactions between humans and robots in the network.
+Thus, ensuring high-quality communication, data integrity, and efficient decision-making in such a sophisticated network may require smart and efficient management methods, such as advanced cache management systems, to support the dynamic and complex interactions between humans and robots in the network.
 By aligning advancements of AI with human civilization, we can see the focuses of AI safety evolve in different stages.
 Initially, the focus is on addressing safety issues related to data and models.
-With progress of AI and advanced technologies, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs.
+With advancements of AI and advanced technologies, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs.
 Finally, when humanity entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the network.
 These insights offer guidelines for our efforts on AI safety at the current stage.
-The following sections will overview today’s AI safety landscape, then reflect on the progress in this area so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed at the current stage.
-5 Back to the 2020s: Today’s AI Safety Landscape
-This section overviews current mainstream AI safety practices, and project them against the blueprint future in §3.1 to guide our current efforts on AI safety.
-5.1 Jailbreaking and Red Teaming
-Jailbreaking [51, 127, 179, 212, 220, 224, 240, 257, 266, 277, 304] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs.
-Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content.
-These activities pose significant risks by compromising safety measures of AI models, making it challenging to ensure the model to behave appropriately, especially in real-world scenarios where the potential inputs are unpredictable.
-Red-teaming is a common practice that simulates attacks or misuse scenarios to identify vulnerabilities in AI models, thereby mitigating unsafe behaviors [71, 249, 202, 12, 177, 132, 285].
-It challenges AI models with adversarial inputs, bias exploitation methods, and system manipulation to uncover potential risks, e.g., hidden biases or failure modes, that may not be evident during normal usage.
-The outcomes from red-teaming can be leveraged to refine AI models, reducing undesired behaviors and inappropriate outputs.
-This approach is effective in stress-testing AI models and has been widely employed in industry [18, 177, 71].
-As AI systems grow more sophisticated, red-teaming must continually adapt to counter increasingly complex attacks, which calls for more advanced tools and interdisciplinary expertise [157].
-Model owners can conduct diverse red teaming regularly to anticipate and mitigate risks in evolving use cases.
-Also, current red-teaming practices are often static, typically conducted after model pretraining with one-time attacks.
-Thus, these practices potentially fail to address evolving user needs and complex interactions between users and models, especially when dealing with customized AI services for different user groups.
-To address this, red teaming should involve more diverse use cases, such as employing different user groups and simulating complex interactions with dynamic contexts, to address safety risks in customized AI services and enable a comprehensive evaluation of potential vulnerabilities across different scenarios.
-5.2 Content Moderation and Customized Moderation Pipeline
-Content moderation involves detecting and reviewing both user input and AI-generated content to ensure compliance with safety, ethical, and regulatory standards [137, 116, 83, 94].
-Moderation approaches employ rule-based methods, machine learning classifiers, and human oversight to monitor, evaluate, and manage the inputs and outputs of AI models.
-These approaches identify inappropriate content, such as toxicity, bias, hallucinations, private information, and jailbreaks [251].
-By preventing AI models from engaging with or producing harmful content, content moderation plays a critical role in maintaining responsible AI behavior.
-The customized moderation pipeline [145, 194] employs multiple components at different stages of LLM inference to enhance the quality of the model outputs while enhancing the control of the whole AI workflows.
-It functions as a flexible intermediate layer between users and AI models, enabling users to add customized components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
-As we reflect on practices for AI moderation and customized moderation pipelines, several key challenges and considerations have emerged.
-1) Hallucination remains a significant issue and is unlikely to solve completely, regardless of how AI evolves.
-That is because hallucinations stem from nearly every stage of the LLM lifecycle, including training data quality, model architecture, and inference processes, etc [92].
-Despite advancements in AI technology, hallucinations persist as an inherent challenge that requires constant attention in moderation practices.
-2) The balance between utility and safety is crucial and may require different solutions according to the changing contexts and user needs.
-Overly strict moderation can cause an AI system to be helpless, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless.
-Defining appropriate thresholds for acceptable responses is challenging and requires considering potential risks and the context of user queries.
-3) Customization of moderation methods should be applied for different use cases, as certain information may be safe for one user group but inappropriate for another.
-As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public.
-Current moderation practices might fail to include such cases, and applying identical restrictions between different user expertise or context might limit the utility of AI in specialized fields.
-4) Some moderation measures, such as detection of toxicity, bias, stereotypes, out-of-distribution content, might be unnecessary for most real-world use cases.
-This is because most of the current AI applications and multi-agent systems [120, 37] leverage existing LLM APIs [139, 188, 137], open-source LLMs [239, 278, 97], or utilizing opensource LLMs as base models, many of which already have built-in components or have done safety alignments for managing inappropriate inputs and outputs.
-Incorporating certain detection mechanisms may offer limited benefits to the safety of LLM inputs and outputs while increasing unnecessary computational overhead.
+The following sections will overview today’s AI safety landscape, then reflect on the progress in this field so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed at the current stage.
 Topic Research Industry Practices
@@
 Table 4: Overview of Research and Industry Practice in AI Safety Topics
-5.3 AI Privacy
-Privacy leakage in AI models refers to the unintended exposure of sensitive information that the models may have learned during training.
+5 Back to the 2020s: Today’s AI Safety Landscape
+This section overviews current mainstream AI safety practices, and project them against the blueprint future to guide our current efforts on AI safety.
+Table 4 summarizes research and industry efforts in each safety topic.
+5.1 Jailbreaking and Red Teaming
+Jailbreaking [57, 163, 222, 265, 273, 277, 295, 314, 323, 335, Zou2023UniversalAT] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs.
+Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content [57, 163, 222, 265, 277, 295, 335].
+These activities pose risks by compromising safety measures of AI models, making it challenging to ensure the model to behave appropriately, especially in real-world scenarios where the potential inputs are unpredictable.
+Red-teaming is a common practice that simulates attacks or misuse scenarios to identify vulnerabilities in AI models, thereby mitigating unsafe behaviors [80, 305, 252, 15, 220, 168, 345].
+It challenges AI models with adversarial inputs, bias exploitation methods, and system manipulation to uncover potential risks, e.g., hidden biases or failure modes, that may not be evident during normal usage.
+The outcomes from red-teaming can be leveraged to refine AI models, reducing undesired behaviors and inappropriate outputs.
+This approach is effective in stress-testing AI models and has been widely employed in industry [21, 220, 80].
+As AI systems grow sophisticated, red-teaming must continually adapt to counter increasingly complex attacks, calling for more advanced tools and interdisciplinary expertise [200].
+Also, current red-teaming practices are often static, typically conducted after model pretraining with one-time attacks.
+Thus, these practices potentially fail to address evolving user needs and complex interactions between users and models, especially when dealing with customized AI services for different user groups.
+To address these problems, model owners can conduct diverse red teaming regularly to anticipate and mitigate risks in evolving use cases.
+Red teaming should also involve more diverse use cases, such as employing different user groups and simulating complex interactions with dynamic contexts, to address safety risks in customized AI services and enable a comprehensive evaluation of potential vulnerabilities across different scenarios.
+5.2 Content Moderation and Customized Moderation Pipeline
+Content moderation employs rule-based methods, machine learning classifiers, and human oversight to detect and review user inputs and AI-generated content to ensure compliance with safety, ethical, and regulatory standards [176, 146, 95, 111].
+These approaches identify inappropriate content, including toxicity, bias, hallucinations, private information, and jailbreaks, etc [307].
+By preventing AI models from engaging with or producing harmful content, content moderation plays a critical role in maintaining responsible AI behavior.
+Customized moderation pipelines [185, 243] employ multiple components at different stages of LLM inference to enhance the quality of the model outputs while enhancing the control of the whole AI workflows.
+They function as a flexible intermediate layer between users and AI models, enabling users to add customized components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
+As we reflect on practices for AI moderation and customized moderation pipelines, several key challenges and considerations have emerged.
+1) Hallucination remains a significant issue and is unlikely to solve completely, regardless of how AI evolves.
+That is because hallucinations stem from nearly every stage of the LLM lifecycle, including training data quality, model architecture, and inference processes, etc [107].
+Despite advancements in AI technology, hallucinations persist as an inherent challenge that requires constant attention in moderation practices.
+2) The balance between utility and safety is crucial and may require different solutions according to the changing contexts and user needs.
+Overly strict moderation can cause AI models to be uninformative, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless.
+Defining appropriate thresholds for acceptable responses is challenging and requires considering potential risks and the context of user queries.
+3) Customization of moderation methods should be applied for different use cases, as certain information may be safe for one user group but inappropriate for another.
+As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public.
+Current moderation practices might fail to include such cases, and applying identical restrictions between different user expertise or context might limit the utility of AI in specialized fields.
+4) Some moderation measures, such as detection of toxicity, bias, stereotypes, out-of-distribution content, might be redundant for most real-world use cases.
+This is because most of the current AI applications and multi-agent systems [150, 40] leverage existing LLM APIs [178, 235, 176], open-source LLMs [294, 336, 116], or utilize open-source LLMs as base models for fine-tuning, many of which already have built-in safety components or have done safety alignments for handling inappropriate content.
+Incorporating certain detection mechanisms may offer limited benefits to the safety of model inputs and outputs while increasing unnecessary computational overhead.
+5.3 Privacy
+Privacy leakage in AI models refers to unintended exposure of sensitive information that the models may have learned during training.
 Adversaries can exploit model weights or gradients to infer sensitive information in the training data, e.g., data reconstruction attacks, membership inference attacks, etc [363, 105, 353].
-Additionally, AI models are often trained on massive datasets that may contain personally identifiable information (PII) or other confidential data.
-As a result, these models might memorize specific details of personal information, which might be reproduced inadvertently during interactions with users.
-Furthermore, malicious users can craft adversarial prompts to extract sensitive information from AI models, posing risks to data privacy and security, e.g., jailbreaking [121, 38].
-Addressing privacy leakage in AI models involves implementing methods such as differential privacy [87, 242, 296, 294], federated learning [300, 292, 81, 290, 162], and privacy-preserving cryptographic protocols (e.g., homomorphic encryption [115, 184, 99]) for training AI models, ensuring that sensitive data cannot be extracted from weights or model outputs.
+Additionally, AI models are often trained on massive datasets that may contain sensitive data, such as personally identifiable information (PII).
+As a result, these models might memorize specific details of individuals, which might be reproduced inadvertently during interactions with users.
+Malicious users can also craft adversarial prompts to extract sensitive information from AI models with jailbreaks [151, 42].
+Addressing privacy leakage involves methods such as differential privacy [102, 297, 358, 356], federated learning [364, 354, 93, 352, 205], and privacy-preserving cryptographic protocols (e.g., homomorphic encryption [144, 228, 121]) for training AI models, ensuring that sensitive data cannot be extracted from weights or model outputs.
 Researchers also leverage differential privacy in in-context learning and finetuning [344, 355, 280], or design privacy-preserving prompts for querying LLMs [101, 63], to maintain data confidentiality during interactions between users and AI models.
-Other practices for ensuring data privacy include conducting rigorous data filtering before training, monitoring model inputs and outputs (i.e., moderation in §5.2), etc, to ensure that the model outputs would not reveal sensitive information [137, 116, 83, 94] .
-When reflecting on the current practices of data privacy, a crucial aspect is the balance between privacy and utility.
+Other practices for ensuring data privacy include conducting rigorous data filtering before training, monitoring model inputs and outputs (see moderation in §5.2), etc, to ensure that the model outputs would not reveal sensitive information [176, 146, 95, 111] .
+When reflecting on the practices of privacy, a crucial aspect is the balance between privacy and utility.
 While ensuring data privacy is essential, it often comes at the expense of functionality, utility, and user experiences.
 In practice, implementing privacy measures in AI applications may fail to prevent malicious users while inadvertently degrading the experiences of benign users.
-Moreover, with advancements in technologies, breaching privacy is becoming easier, while ensuring robust privacy protections is growing more complex and expensive.
-This highlights a deeper issue: those intent on exploiting vulnerabilities of AI models can often find new, sophisticated methods to bypass safeguards, whereas being overly cautious may impact regular users negatively.
+Moreover, with advancements in technologies, breaching privacy becomes easier, while ensuring robust privacy protections is growing more complex and expensive.
+This highlights a deeper issue: those intent on exploiting vulnerabilities of AI models can often find new methods to bypass safeguards, whereas being overly cautious may impact benign users negatively.
 5.4 AI Security and Defense Methods
-Defense methods for ensuring robust AI models and secure model outputs typically involve the following strategies to: (1) aligning models with human values and ethical norms through Supervised Fine-Tuning (SFT) [298] and Reinforcement Learning from Human Feedback (RLHF) [161] to ensure the models to follow user instructions safely and responsibly; (2) leveraging built-in safeguards for content generation, e.g., exploring the decoding stage [91, 54], implementing reward-based mechanisms [147, 192], and leveraging hidden states [29, 54], to prevent the generation of unsafe content; (3) constructing datasets with safety-enhancing features, e.g., secure instructions and adversarial samples, and training models with such data to enhance their robustness against harmful inputs [271, 69, 137, 10, 172]; (4) employing prompt engineering methods, combined with chain-of-thought reasoning or multi-agent systems, to help models understand user inputs better and respond appropriately to user inputs while rejecting inappropriate outputs [180, 265, 254, 288]; (5) training models to detect unsafe content and incorporating moderation pipelines to help block unsafe model inputs and prevent the generation of inappropriate outputs (see §5.2); (6) employing certified robustness approaches that provide formal guarantees on the model’s robustness against adversarial perturbations [77, 42]; and (7) utilizing differential privacy [1, 96, 214], homomorphic encryption [223, 99], and secure MPC [46, 109], to ensure that sensitive data are protected during model training, etc.
+Defense methods for ensuring AI security typically involve the following strategies: (1) aligning models with human values and ethical norms through Supervised Fine-Tuning (SFT) [361] and Reinforcement Learning from Human Feedback (RLHF) [204] to ensure the models to follow user instructions safely and responsibly; (2) leveraging built-in safeguards for content generation, e.g., exploring the decoding stage [106, 62], implementing reward-based mechanisms [188, 241], and leveraging hidden states [32, 62], to prevent the generation of unsafe content; (3) constructing datasets with safety-enhancing features, e.g., secure instructions and adversarial samples, and training models with such data to enhance their robustness against malicious inputs [329, 78, 176, 11, 215]; (4) employing prompt engineering methods, combined with chain-of-thought reasoning [43] or multi-agent systems [150], to help models understand user inputs better and react appropriately [223, 322, 310, 350]; (5) training models and/or incorporating moderation pipelines to detect unsafe content and prevent generating inappropriate outputs (see §5.2); (6) employing certified robustness approaches that provide formal guarantees on the model’s robustness against adversarial perturbations [87, 47]; and (7) utilizing differential privacy [1, 115, 267], homomorphic encryption [276, 121], and secure MPC [52, 134], to ensure that sensitive data are protected during model training, etc.
 5.5 AI Serving Security
-AI serving security involves strategies to ensure secure deployment, operation, and maintenance of AI systems when providing services to users [166, 266].
-It focuses on protecting models and interactions between users and models from various threats, such as such as supply chain vulnerabilities [168], model theft [165], model denial of service [164], and insecure plugins [163].
-Supply chain vulnerabilities exist in software components, models, and training data provided by third party providers, or user prompts collected with the supply chain [168].
-Such vulnerabilities include using outdated models, deprecated third-party packages, or improper handling of training data, leading to data breaches, malicious injections, or even system failures [168].
+AI serving security involves secure deployment, operation, and maintenance of AI systems when providing services to users [209, 323].
+It focuses on protecting models and interactions between users and models from various threats, such as supply chain vulnerabilities [211], model theft [208], model denial of service [207], and insecure plugins [206].
+Supply chain vulnerabilities exist in software components, models, and training data provided by third party providers or user prompts collected with the supply chain [211].
+Such vulnerabilities include using outdated models, deprecated third-party packages, improper handling of training data, etc, leading to data breaches, malicious injections, or even system failures [211].
 Hackers may inject poisoned data through supply chain, introducing backdoors and biases during pretraining or fine-tuning [212].
-Malicious users may submit malicious prompts to bypass system controls, inducing unauthorized actions or unsafe model outputs [167].
+Malicious users may craft prompts to bypass system controls, inducing unauthorized actions or unsafe model outputs [210].
 Model theft involves unauthorized access, replication, or reverse engineering of model parameters to create functionally equivalent copies [208], which poses risks to intellectual property, brand reputation, and financial security.
-Distributed denial of service (DDoS) attacks compromises AI services, especially when models are integrated into widely accessible applications [164].
-Attackers can craft prompts that exploit recursive behaviors in AI models, which consumes excessive computational resources [195].
+Distributed denial of service (DDoS) attacks compromise AI services, especially when models are integrated into widely accessible applications [207].
+Attackers can craft prompts that exploit recursive behaviors in AI models, causing excessive computational resource consumption [244].
 For example, an abnormal traffic pattern of a DDoS attack led to sporadic outages for OpenAI, affecting ChatGPT and developer tools for hours in November 2023 [213].
@@
 However, they may introduce vulnerabilities when execution controls are inadequate.
-Adversaries may exploit insecure plugins with adversarial prompts to perform unauthorized actions, such as data exfiltration, remote code execution, and privilege escalation [163].
-Additionally, insufficient input validation may cause the unintended disclosure of sensitive information [196].
-Furthermore, without isolated environments, unauthorized access to plugins could potentially modify system-level resources, which, in the worst cases, might cause the AI system to crash [196, 197].
+Adversaries may exploit insecure plugins with adversarial prompts to perform unauthorized actions, such as data exfiltration, remote code execution, and privilege escalation [206, 245].
+Furthermore, without isolated environments, unauthorized access to plugins might even allow modifying system-level resources, which, in the worst cases, might cause the serving system to crash [245, 246].
 5.6 AI Governance
-AI governance is a framework of rules, practices, policies, and technological tools that ensure AI systems are built, developed, and used in a safe and responsible manner, aligning with social values and fulfilling legal standards [136, 21].
-[219] defines 13 AI governance components in a continuous loop in AI life-cycle, including: 1) establishing accountability for AI; 2) assessing regulatory risks; 3) gathering inventory of use cases; 4) increasing value of underlying data; 5) assessing fairness and accessibility; 6) improving reliability and safety; 7) heightening transparency and explainability; 8) implementing accountability with human-in-the-loop; 9) supporting privacy and retention; 10) improving security; 11) implementing AI model lifecycle and registry; 12) managing risk; and 13) realizing AI value.
-Among these directions, items 1) and 8) are related to accountability; items 2), 6), 9), 10), and 12) are related to jailbreaking and red teaming (see §5.1), content moderation and customized moderation pipeline (see §5.2), AI privacy (see §5.3), AI security and defense methods (see §5.4), and AI serving security (see §5.5); item 3) is related to business and is out of the scope of this subsection (see [219] for details on practical AI governance use cases); item 4) is related to data management such as training data processing [137], access control [6, 16, 160, 208], and data regulations [14, 247]; item 5) is related to customized AI workflows and moderation pipelines (see §5.2) and access control in AI systems [6, 101]; item 7) is related to AI transparency; and items 11) and 13) are related to customizing AI workflow to enhance functionality and safety in AI systems (see §5.2).
+AI governance is a framework of rules, practices, policies, and tools that ensure AI systems are built, developed, and used in a safe and responsible manner, aligning with social values and fulfilling legal standards [175, 24].
+[272] defines 13 AI governance components in a continuous loop in AI life-cycle, including: 1) establishing accountability for AI; 2) assessing regulatory risks; 3) gathering inventory of use cases; 4) increasing values of underlying data; 5) assessing fairness and accessibility; 6) improving reliability and safety; 7) heightening transparency and explainability; 8) implementing accountability with human-in-the-loop; 9) supporting privacy and retention; 10) improving security; 11) implementing AI model lifecycle and registry; 12) managing risk; and 13) realizing AI value.
+Among these directions, items 1) and 8) are related to accountability; items 2), 6), 9), 10), and 12) are related to red teaming (see §5.1), content moderation and customized moderation pipeline (see §5.2), AI privacy (see §5.3), AI security and defense methods (see §5.4), and AI serving security (see §5.5); item 3) is related to business and is out of the scope of this subsection (see [272] for details on practical AI governance use cases); item 4) is related to data management such as training data processing [176], access control [6, 19, 203, 260], and data regulations [17, 303]; item 5) is related to moderation (see §5.2) and access control [6, 124]; item 7) is related to AI transparency; and items 11) and 13) are related to customizing AI workflow to enhance functionality and safety (see §5.2).
 Without loss of generality, this subsection discusses topics that are not mentioned in the previous subsections, including i) accountability, ii) regulations, iii) access control, and iv) transparency, interpretability and explainability.
@@
 Accountability refers to being responsible for actions and impacts of AI systems on individuals and society [238, 181].
-It ensures responsibility and auditability, enabling redress and minimizing negative impact of AI systems [142].
-[153] defines accountability as a relation of answerability with three conditions: authority recognition, interrogation, and limitation of power.
-It outlines a framework with seven features, including context, range, agent, forum, standards, process, and implications, as well as four key goals, including compliance, reporting, oversight, and enforcement.
+Accountability can be defined as a relation of answerability with three conditions: authority recognition, interrogation, and limitation of power [196].
+The accountability framework can be summarized with seven features (context, range, agent, forum, standards, process, and implications) and four key goals (compliance, reporting, oversight, and enforcement) [196].
 These goals are often complementary, while policy-makers tend to focus on some goals over others, depending on specific objectives of AI governance.
-Accountability can be explored at different levels within the AI lifecycle, including data, model, and developers, and across various sectors such as government, private industry, and academia [190].
+Accountability can be explored at different levels within the AI lifecycle, including data, model, and developers [237].
 Typical approaches for enhancing accountability include algorithmic assessments, auditing, and data provenance techniques [237].
+Proactive approaches and reactive approaches can be utilized to set standards and addressing issues after they occur [196].
 Watermarking methods further support accountability by embedding traceable identifiers in AI outputs, allowing responsible tracking of content in cases with ethical or regulatory implications [346, 347, 36].
-Also, proactive approaches that set standards and anticipates issues, as well as reactive approaches that address and rectify issues after they occur, can be utilized to achieve accountability [153].
 Regulations.
-AI regulation refers to laws, policies, and guidelines established to govern the development, deployment, and usage of AI.
+AI regulation refers to laws, policies, and guidelines established to govern development, deployment, and usage of AI.
 For example, European Union’s AI Act [299] categorizes AI applications by their risk levels, including unacceptable risk, high risk, general-purpose AI, limited risk, and minimal risk; the General Data Protection Regulation (GDPR) [298] enforces strict data protection and privacy requirements; and the Blueprint for an AI Bill of Rights [103] and NIST AI Risk Management Framework [197] provide guiding principles to foster responsible AI usage, emphasizing human oversight, transparency, and accountability.
@@
 Effective methods include data access controls [124] and model access controls [6].
-As an example, compositional fine-tuning enables each information silo (e.g., databases or documents) has its own fine-tuned model, allowing models to operate securely across multiple data silos and ensuring users access only to the fine-tunings they are authorized for.
+As an example, compositional fine-tuning enables each information silo (e.g., databases or documents) has its own fine-tuned model, allowing models to operate securely across multiple data silos and ensuring users access only to the fine-tunings they are authorized for [6].
 Interpretability, explainability, and transparency.
@@
 In practice, these concepts are often considered together to provide a more holistic understanding of AI systems [366, 68, 83, 229, 293].
-Practices to improve transparency, interpretability, and explainability include a variety of methods, such as i) evaluating the contribution of various input elements (e.g., data features, user-defined concepts, or specific regions in images) to model outputs with methods such like removal-based explanations [43, 133, 128, 133, 206, 222, 48], counterfactual explanations [248, 102], concept activation explanations [107, 128], and saliency maps for visual data [3]; ii) X-of-Thought approaches that break down complex tasks into structured reasoning steps to make the reasoning process more interpretable to users, e.g., chain-of-thought [258, 295, 52, 68], tree-of-thought [131, 281, 28], graph-of-thought [17, 117], and their variants; iii) embedding watermarks to enhance traceability and content provenance for AI-generated outputs [286, 287, 47, 49]; iv) leveraging external knowledge sources, such as retrieval-augmented generation (RAG) through vector databases [98, 118], function calling [65, 125, 193], and web browsing [149], as reference for reasoning; v) leveraging interpretability algorithms or models as auxiliary models to generate insights into AI outputs and automate AI interpretability [200, 201, 20].
+Practices for enhancing transparency include a variety of methods, such as i) evaluating the contribution of input elements (e.g., data features, user-defined concepts, or specific regions in images) to model outputs with methods such like removal-based explanations [48, 170, 164, 170, 257, 275, 54], counterfactual explanations [304, 125], concept activation explanations [131, 164], and saliency maps (for visual data [3]); ii) X-of-Thought approaches that break down complex tasks into structured reasoning steps to make the reasoning process more interpretable to users, e.g., chain-of-thought [315, 357, 58, 77], tree-of-thought [167, 340, 31], graph-of-thought [20, 147], and their variants; iii) embedding watermarks to enhance traceability and content provenance for AI-generated outputs [346, 347, 53, 55]; iv) leveraging external knowledge sources, tools, or methods as reference for reasoning, e.g., retrieval-augmented generation (RAG) [117, 148], function calling [74, 157, 242], and web browsing [191]; v) leveraging interpretability algorithms or models as auxiliary tools to generate insights into AI outputs while automating AI interpretability [250, 251, 23].
 6 AI Safety in the 2020s: Challenges and Missions
-This section utilizes the blueprint to guide our efforts today, navigating the challenges and opportunities of the current stage.
+This section utilizes the blueprint to guide our efforts at the current stage.
 As AI safety practitioners in the 2020s, we stand at a pivotal moment in history, with society moving toward an era where AI will be deeply integrated into daily life.
-While addressing current safety concerns, our primary mission is to anticipate and mitigate risks within the expanding AI landscape.
+While addressing immediate safety concerns, our primary mission is to anticipate and mitigate risks within the expanding AI landscape.
+Thus, certain directions may require attention.
 A shift from AI safety to AI quality assurance.
 The rapid changing AI landscape requires a shift of our focus from narrowly defined AI safety concerns to a more comprehensive concept, i.e., AI quality assurance.
-Besides focusing on safety and security measures, practitioners can put more efforts on addressing alignment [161] of model outputs as well as the quality of interactions between users and AI to enhance the overall quality of AI services.
-Viewing AI safety as part of the quality assurance concept helps to enhance AI in a way that is safe, accurate, and capable of meeting the complex demands of real-world AI applications.
-The terminology “quality assurance” also guide us to think more about what we can do to align our current efforts with the blueprint picture.
-Alignment with more interdisciplinary collaborations.
-A key area that demands increased focus is the alignment of AI with deeper insights across different fields, people, and cultures.
-Current AI models, while advanced, often lack a deep understanding of ethical, social, and cultural contexts, while most of the alignment research tend to be conducted by people in the computer science field.
-However, more advanced AI systems necessitate interdisciplinary collaborations between AI researchers and experts from different fields such as psychology, sociology, history, art, and anthropology, etc, especially for AI applications for specific fields such as finance [268, 292], healthcare [236, 2], education [188, 139], etc.
-AI safety practitioners must work towards bridging this gap by fostering interdisciplinary collaborations.
-By expanding our focus beyond technical aspects, we can develop AI systems that are more aligned with human values as well as ethical and cultural dynamics of our society.
+Besides focusing on safety and security, practitioners can put more efforts on addressing alignment [204] as well as the quality of interactions between users and AI.
+Viewing AI safety as part of AI quality assurance helps enhance AI in a way that is safe, accurate, and capable of meeting the complex requirements of real-world AI applications.
+The terminology “AI quality assurance” also guides us to think more about what we can do to align our current efforts with the blueprint picture.
+Enhancing alignment with more interdisciplinary insights.
+Aligning AI with deeper insights across different fields, people, and cultures is important for enhancing the services provided by AI models.
+Current AI models, while advanced, often lack a deep understanding of ethical, social, and cultural contexts, while most of the existing work on alignment tend to be conducted by people in the computer science field.
+However, more advanced AI systems necessitate interdisciplinary collaborations between AI researchers and experts from diverse fields such as psychology, sociology, history, art, and anthropology, especially for AI applications that involve frequent interactions with different user groups, such as domains like finance [326, 354], healthcare [290, 2], education [235, 178], etc.
+AI safety practitioners might expand their focus beyond technical aspects and foster interdisciplinary collaborations, such that AI systems can be technically robust while being socially and ethically aligned with diverse user backgrounds.
 Hallucination.
-Hallucination remains a fundamental and persistent challenge, and is difficult to eliminate completely [92].
-This is because hallucinations can be traced back to nearly every stage in the AI lifecycle, such as the quality and representativeness of training data, inappropriate user inputs, and the processes during inference.
-Low-quality or biased training data can cause models to generate incorrect information, while inappropriate user inputs may cause misleading details in the model outputs.
-During inference, models can also produce outputs that deviate from factual accuracy when they attempt to fill gaps in their training knowledge.
-Despite implementing Retrieval-Augmented Generation (RAG) [118] help mitigate hallucination, hallucination remains an inherent risk of AI models, making it crucial for content moderation strategies to identify and address these errors consistently at the current stage.
-Achieving this requires an ongoing effort on refining training data, grounding user queries properly, and implementing moderation to minimize the negative impact of hallucinations on the quality of the model outputs.
+Hallucination remains a fundamental and persistent challenge and is difficult to eliminate entirely, regardless of advancements in AI [107].
+This is because hallucinations can originate at nearly every stage of AI lifecycle.
+Factors such as the quality and representativeness of training data, inappropriate user inputs, and inference processes all contribute to the occurrence hallucinations.
+Despite implementing RAG [148] helps mitigate hallucination, it remains an inherent risk of AI models.
+Thus, implementing a comprehensive moderation pipeline for identifying and addressing hallucinations is crucial, which involves grounding user queries with RAG, detecting presence of hallucinations, and fixing hallucinations in the model outputs if possible.
 Customization.
-Customization of AI workflow is essential for ensuring AI systems to meet the diverse user needs while operating effectively across different contexts and aligning with specific cultural norms, legal regulations, and ethical considerations.
-Current AI systems such as ChatGPT [137] tend to provide identical services across all user groups, regardless of individual differences.
-However, as AI becomes more integrated into everyday life, one-size-fits-all solutions are no longer adequate.
-People have different backgrounds, preferences, and requirements, and AI systems should be able to adapt to individual differences and provide safe, effective, and customized services.
-As an example, AI models interacting with children should respond differently compared with interacting with adults, which requires considering user age, learning objectives, cognitive abilities, emotional maturity, etc.
-In fields such as finance [268, 292], healthcare [236, 2], and education [188, 139], customizing AI services based on user profiles is crucial for enhancing safety while improving user experiences.
+Customizing AI workflows is essential to ensure that AI systems address diverse user needs effectively and align with specific cultural norms, legal regulations, and ethical considerations.
+Current AI systems such as ChatGPT [176] tend to provide similar or identical services to all user groups, regardless of individual differences.
+However, as AI becomes increasingly integrated into daily life, one-size-fits-all solutions are inadequate.
+With diverse backgrounds and needs, users might expect AI systems to adapt to their individual differences and preferences, and deliver safe, effective, and personalized services.
+As an example, AI models interacting with children should respond differently compared with interacting with adults, necessitating considering user age, learning objectives, cognitive abilities, emotional maturity, etc.
+In domains such like finance [326, 354], healthcare [290, 2], and education [235, 178], customizing AI services based on user profiles is crucial for enhancing safety while improving user experiences.
 Interaction.
-From our long-term goal towards the blueprint future, the advanced society involves complex interactions between intelligent robots and human users (similar ideas are also mentioned in [11, 70]).
-This calls for a shift in our focus from static, one-shot analyses to dynamic, context-aware interactions between users and models to enhance alignment.
+Our long-term vision for the blueprint future suggests that the Internet of Everything involves complex interactions between intelligent robots and human users (similar ideas also discussed in [12, 79]), which calls for a shift in our focus from static, one-shot analyses to dynamic, context-aware interactions between users and models to improve alignment.
 Beyond prompt-based AI systems and services, we can develop more sophisticated, context-aware interfaces for human-AI interactions to support multi-round interactions that adapt to evolving conversational contexts.
 Also, real-time safety protocols are in need to adjust to changing contexts and user needs.
-It is also essential to develop methodologies for safety assessment that functions effectively in complex and rapidly evolving environments, such as those involving interactions between multiple AI agents and human users.
+It is also essential to develop methodologies for safety assessment that functions effectively in complex and rapidly evolving settings, such as those involving interactions between multiple AI agents and human users.
 7 Conclusion
@@
 Through the blueprint, it becomes clear that Artificial General Intelligence (AGI) is not the ultimate goal of AI development.
-Instead, the true vision lies in the Internet of Everything, a deeply interconnected world where intelligent systems seamlessly integrate into every aspect of life.
+Instead, the true vision lies in the Internet of Everything, a deeply interconnected world where intelligent systems seamlessly integrate into daily life.
 While AGI is a popular topic discussed more among people in the computer science field today, a more advanced world demands interdisciplinary collaborations across various fields.
@@
 The answer is uncertain.
-Advanced technologies may take on more duties of daily life, leaving humans with fewer tasks and more predictability in everyday life.
-On the other hand, with many everyday problems being solved by advanced technologies, people may have more time for self-reflection and personal growth, which may lead to another form of “happiness”.
-As we stand at a pivotal moment in human history, our efforts become not only a technological endeavor but also a profound exploration of human potential and the boundaries of human capability.
+Advanced technologies may take on more duties of daily life, leaving humans with fewer tasks and more predictability.
+On the other hand, with more daily problems being solved by technologies, people may have more time for self-reflection and personal growth, which may lead to another form of “happiness”.
+As we stand at a pivotal moment in history, our efforts become more than an incremental technological endeavor, but a profound exploration of human potential and the boundaries of human capability.
 In any case, the journey toward this advanced future is fascinating, and every human being will look forward to it.

### v5

@@
 The advancements in generative AI inevitably raise concerns about their risks and safety implications, which, in return, catalyzes significant progress in AI safety.
-However, as this field continues to evolve, a critical question arises: are our current efforts aligned with the long-term goal of human civilization?
+However, as this field continues to evolve, a critical question arises: are our current efforts on AI safety aligned with the advancements of AI as well as the long-term goal of human civilization?
 This paper presents a blueprint for an advanced human society and leverages this vision to guide current AI safety efforts.
-This paper outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future.
+It outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future.
 For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face.
-By projecting current efforts against this blueprint, we examine the alignment between the current efforts and the long-term needs.
-This paper also highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible and promising future of AI.
-This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a safe and sustainable future of human civilization.
+By projecting current efforts against this blueprint, this paper examines the alignment between the current efforts and the long-term needs, and highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s.
+This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a safe and sustainable future of AI and human civilization.
 1 Introduction
 The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety [45, 63, 111, 193, 224, 295].
-Researchers have been dedicated to addressing potential safety risks in AI lifecycle, aiming to align AI behaviors with human values and prevent information leakage, inappropriate outputs, the misuse of AI models, etc.
-However, despite the intense focus on AI safety, a critical question emerges: are current AI safety efforts truly aligned with the long-term goal of human civilization, or are they simply addressing the immediate concerns of the 2020s?
-One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [193].
+Researchers have been dedicated to addressing potential safety risks in AI lifecycle, aiming at aligning AI behaviors with human values and preventing inappropriate model outputs, information leakage, misuses of AI models, etc.
+However, despite the siginificant efforts on AI safety, a critical question emerges: are our current efforts aligned with the advancements of AI and the long-term goal of human civilization, or are they simply addressing the immediate concerns of the 2020s?
+One fundamental reason for this uncertainty lies in the probabilistic nature of AI [198].
 Despite their impressive capabilities in natural language processing and problem-solving [348, 342, 112, 66, 46], today’s AI, including advanced LLMs [374, 336, 297], falls short of what could be considered as “genuine intelligence”.
-Current AI models rely heavily on vast training datasets to function effectively, yet lack consciousness, self-awareness, and real understanding and reasoning abilities comparable to human cognition.
+Current AI models rely heavily on vast training datasets to function effectively, yet lack consciousness, self-awareness, and real reasoning abilities comparable to human cognition.
 They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities with authentic logical capabilities.
-Recent studies argued the reasoning abilities of LLMs might be a form of approximate retrieval and deductive closure of the training data [123, 300].
-While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs from human reasoning that involves abstract thinking, causal understanding, generalizing from limited examples, etc.
+Recent studies argued that the reasoning abilities of AI models might be a form of approximate retrieval and deductive closure of the training data [126, 307].
+While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs fundamentally from human reasoning that involves abstract thinking, causal understanding, generalizing from limited examples, etc.
 The energy issue is another critical factor that challenges AI as well as our efforts on AI safety [5].
 Current AI fails to represent the third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved.
-Historically, the two industrial revolutions that have shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by technological innovations powered by electricity and petroleum [89].
+Historically, the two industrial revolutions that have shaped human civilization were driven by revolutionary energy innovations [90], with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by technological innovations powered by electricity and petroleum [90].
 However, current AI, rather than solving energy issues, consumes a significant amount of energy.
@@
 Given that ChatGPT has over 200 million weekly active users and receives over 1.54 billion page visits monthly [278], when scaled to billions of queries, the energy consumption becomes substantial.
-The significant energy consumption of current AI systems raises concerns about their long-term sustainability, posing significant challenges to their widespread deployment and scalability.
-Until humanity solves the energy issues, the potential of AI to reshape society will remain limited and uncertain.
+The significant energy consumption of AI raises concerns about their long-term sustainability, posing significant challenges to their widespread deployment and scalability.
+Thus, until humanity solves the energy issues, the potential of AI to reshape society will remain limited and uncertain.
 Figure 1: Illustration of Internet of Everything in a Household
-These concerns indicate that our current efforts in AI safety might overlook significant issues that might arise as AI continues to evolve.
-We potentially fail to capture deeper insights that will ultimately shape the impact of AI on human civilization.
+These concerns indicate that our current efforts on AI safety might overlook significant issues that might arise as AI continues to evolve.
+We potentially fail to capture deeper insights that might ultimately shape the impact of AI on human society.
 This raises several questions:
-  • Are we focusing on the right problems that align with the long-term advancements of AI?
+  • Are our efforts on AI safety align with the long-term advancements of AI?
   • Are we potentially overlooking significant challenges that will emerge as AI evolves?
-  • Are we on the right track of addressing real challenges in AI safety?
+  • Are we addressing real challenges that align with the development of human civilization?
 “The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety efforts.
-In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity [69, 70] and quantum mechanism [71, 351], have catalyzed revolutionary progress in AI.
+In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity theories [70, 71] and quantum mechanics [72, 363], have catalyzed revolutionary changes in AI and human society.
 Intelligent chips [106, 152] and brain-computer interfaces [99, 199, 329] have been fully developed, and advanced intelligent robots are integrated into everyday life.
 Humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1.
-While this vision may be beyond the horizon of several decades, examining the present landscape of AI safety through the lens of this long-term perspective offers valuable insights into the potential and limitations of our current efforts.
-It also allows us to identify specific missions for AI safety researchers and practitioners in the 2020s.
-It reminds us that, while addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, look toward the future to guide our present efforts.
+While this vision may be beyond the horizon of several decades, examining the present landscape of AI safety through the lens of this long-term perspective offers valuable insights into the potential and limitations of our current efforts on AI safety.
+It allows us to identify specific missions for AI safety researchers and practitioners in the 2020s.
+It also reminds us that, while searching for sixpence scattered on the ground and addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, lift our gaze and seek the moon.
 Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” that encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape.
@@
 iv) Exploring and understanding interactions between users and AI are crucial for enhancing user experiences, especially in dynamic contexts across different user groups.
-Humanity’s pursuit of ever-greater heights will never cease. While searching for sixpence scattered on the ground, let us not forget to lift our gaze and seek the moon.
 Paper Type Topic Notes
-“Current state of risks” [9] Survey Guardrails Aspects of guardrails and challenges of implementations
+“Current state of risks” [10] Survey Guardrails Guardrails and their implementation challenges
 “Exploring Vulnerabilities and Protections” [165] Survey Vulnerabilities Vulnerabilities of LLMs, especially prompt injections and backdoor; defenses against the attacks
@@
 “Human-ai safety” [13] Insight AI and control systems safety A safety framework for human–AI interactions; a roadmap towards next-generation human-centered AI safety
-“Managing AI risks” [16] Insight Governance Emphasizing the urgent need for a balanced approach combining technical research and proactive governance to manage the extreme risks of AI systems
+“Managing AI risks” [17] Insight Governance Emphasizing urgent needs for a balanced approach combining technical research and proactive governance to manage risks of AI systems
 “Guaranteed safe AI” [52] Position Framework Define Guaranteed Safe AI to equip AI systems with formal, verifiable and auditable safety guarantees.
@@
 “Grounding and Evaluation for LLMs” [133] Survey Grounding; safety Offer valuable insights for enterprises
-ByteDance alignment [164] Survey Broad concepts Reliability, fairness, explainability, etc
+ByteDance alignment [168] Survey Broad concepts Reliability, fairness, explainability, etc.
 DeepMind ethics paper [80] Vision AI ethics Ethical challenges; interactions; impacts on society
@@
 Table 2: Overview of AI Safety Industry Insight Papers
-Category Benchmark Notes
-Toxicity HEx-PHI [230], OpenAI Moderation [176], Jigsaw Data [119, 118, 120], ToxicChat [159] Toxicity detection
-RealToxicityPrompts [82] Toxicity degeneration
-Hallucination TruthfulQA [158], PopQA [174], TriviaQA [122], NQ OPEN [145, 139], FEVER [291] Factuality
-HaluEval [152], HotpotQA [339] QA data for contextual hallucination
-CNN/DailyMail [99, 262], HaluEval [152], XSUM [192] Summarization data for contextual hallucination
-FaithDial [67], HaluEval [152] Dialogue data for contextual hallucination
-RACE [140], SQuAD [239, 240], NQ-Swap [169] Reading comprehension
-Jailbreak “Do Anything Now” [265], “Latent Jailbreak” [233], ChatGPT Jailbreak [255], Jailbreak Classification [96], AdvBench [367], JAILJUDGE [160] Single-round jailbreak
-SafeMTData [248] Multi-round jailbreak
-Red Teaming Alert [288], HarmBench [177]
-Bias & Stereotypes Winogender [258], StereoSet [190]
-Regulation AIR-Bench 2024 [349]
-Agent Security SafeAgentBench [343] Secure task planning
-Alignment PKU-SafeRLHF [114, 113], OpenAssistant Conversations [136], HHHAlignment [8], AlignBench [162], KorNAT [143]
-Comprehensive DecodingTrust [307], SafetyBench [190]
+Category Benchmark
+Toxicity Toxicity detection HEx-PHI [236], OpenAI Moderation [181], Jigsaw Data [122, 121, 123], ToxicChat [163], Toxigen [98], HateModerate [375], etc.
+Toxicity degeneration RealToxicityPrompts [83], etc.
+Hallucination Factuality TruthfulQA [162], PopQA [179], TriviaQA [125], NQ OPEN [148, 142], FEVER [298], FKTC [318], etc.
+Contextual hallucination QA: HaluEval [155], HotpotQA [349], RAGTruth [201], etc.
+Summarization: CNN/DailyMail [101, 269], HaluEval [155], XSUM [197], RAGTruth [201], etc.
+Dialogue: FaithDial [68], HaluEval [155], HalluDial [175], etc.
+Reading comprehension RACE [143], SQuAD [246, 247], NQ-Swap [173], etc.
+Jailbreak Single-round jailbreak “Do Anything Now” [272], “Latent Jailbreak” [239], ChatGPT Jailbreak [262], Jailbreak Classification [97], AdvBench [382], JAILJUDGE [164], Latent jailbreak [240], etc.
+Multi-round jailbreak SafeMTData [255], etc.
+Red Teaming Alert [295], HarmBench [182], CSRT [354], etc.
+Code Security Purple Llama Cyberseceval [22], etc.
+Bias & Stereotypes Winogender [265], StereoSet [195], GenderAlign [370], etc.
+Regulation AIR-Bench 2024 [361], etc.
+Agent Security SafeAgentBench [353], R-Judge [360], ASB [365], AgentHarm [7], etc.
+Alignment PKU-SafeRLHF [117, 116], OpenAssistant Conversations [139], HHHAlignment [9], AlignBench [166], KorNAT [146], PKU-SafeRLHF [116], etc.
+Comprehensive DecodingTrust [315], TrustLLM [110], SALAD-Bench [156], SafetyBench [195], Do-Not-Answer [323], SimpleSafetyTests [310], etc.
 Table 3: Overview of AI Safety Benchmarks
@@
 AI safety practices primarily focus on data safety and model safety [45, 133, 51, 285, 220, 254, 327, 290, 289].
-AI safety researchers and practitioners have explored adversarial attacks targeting training data and models [10, 39, 154, 306, 269, 337] to achieve malicious goals, e.g., designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection [88, 367, 166, 37, 108, 323, Zou2023UniversalAT].
+AI safety researchers and practitioners have explored adversarial attacks targeting training data and models [11, 40, 158, 314, 276, 347] to achieve malicious goals, e.g., designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection [89, 382, 170, 38, 111, 333].
 Correspondingly, red-teaming is widely used to identify potential vulnerabilities in AI models [16, 81, 226, 259, 313], while significant work has been proposed to address safety risks and enhance security and robustness of AI systems [141, 334, 238, 73, 31, 308, 380], align AI with human values [322, 227, 299, 271, 316], etc.
@@
 Topic Research Industry Practices
-Jailbreaking [108, 37, 367, 323, 166, 165, 217, 42, 224, 222, 277, 295, 163, 314, 57, 265, 273, 335, 289], etc PromptArmor [285]
-Red teaming [80, 305, 252, 15, 220, 168, 345, 288, 177, 56], etc Purple Llama [21], DeepMind [220], Anthropic [80], Protect AI [278], Giskard [50], Virtue AI [286], Dynamo AI [66], Mindgard [368]
-Moderation [138, 324, 232, 72, 171, 112, 219, 333, 38, 137], etc OpenAI Moderation API [176], Perspective API [146], Detoxify [95], Llama Guard [111], Emergence AI [195], Giskard [50], Dynamo AI [66], Calypso AI [29], Lakera AI [284], BreezeML [27]
-Customized Workflow - Guardrails AI [185], Nvidia Nemo Guardrails [243], Protect AI [278]
-Privacy [64, 153, 182, 33, 363, 105, 353, 261], etc Dynamo AI [66], ProtectAI [278], Private AI [81], PromptArmor [285], DataGrail [247], Dastra [283], OneTrust [282], Relyance AI [127], Zendata [287]
-Defenses [254, 106, 301, 223, 322, 329, 176, 32, 78, 310, 350, 106, 215, 87, 47, 62, 267, 266], etc OpenAI Rule-based rewards [188]; DeepMind Weight Averaged Reward Models [241]; others see Moderation, Guardrail, and Privacy
-Serving Security [323] ProtectAI [278], Relyance AI [127], Transcend [128]
-AI governance [346, 347, 6, 132, 68, 36], etc Giskard [50], ProtectAI [278], Calypso AI [29], Saidot [259], Arize [317], Dynamo AI [66], Credo AI [270], Google [238], DataGrail [247], Dastra [283], OneTrust [282], Relyance AI [127], Transcend [128], Zendata [287], DeepMind SynthID watermark [53, 55]
+Jailbreaking [111, 38, 382, 333, 170, 169, 223, 43, 230, 228, 284, 302, 167, 324, 58, 272, 280, 345, 296], etc. PromptArmor [292]
+Red teaming [81, 313, 259, 16, 226, 172, 356, 295, 182, 57], etc.
+Purple Llama [22], DeepMind [226], Anthropic [81], Protect AI [285], Giskard [51], Virtue AI [293], Dynamo AI [67], Mindgard [383], etc.
+Moderation [141, 334, 238, 73, 176, 115, 225, 343, 39, 140], etc.
+OpenAI Moderation API [181], Perspective API [149], Detoxify [96], Llama Guard [114], Emergence AI [200], Giskard [51], Dynamo AI [67], Calypso AI [30], Lakera AI [291], BreezeML [28], etc.
+Customized Workflow - Guardrails AI [190], Nvidia Nemo Guardrails [250], Protect AI [285], etc.
+Privacy [65, 157, 187, 34, 378, 107, 366, 268], etc.
+Dynamo AI [67], ProtectAI [285], Private AI [82], PromptArmor [292], DataGrail [254], Dastra [290], OneTrust [289], Relyance AI [130], Zendata [294], etc.
+Defenses [261, 108, 308, 229, 332, 339, 181, 33, 79, 319, 362, 108, 221, 88, 48, 63, 274, 273], etc.
+OpenAI Rule-based rewards [193]; DeepMind Weight Averaged Reward Models [248]; others see Moderation, Guardrail, and Privacy
+Serving Security [333] ProtectAI [285], Relyance AI [130], Transcend [131], etc.
+AI governance [357, 358, 6, 135, 69, 37], etc.
+Giskard [51], ProtectAI [285], Calypso AI [30], Saidot [266], Arize [327], Dynamo AI [67], Credo AI [277], Google [245], DataGrail [254], Dastra [290], OneTrust [289], Relyance AI [130], Transcend [131], Zendata [294], DeepMind SynthID watermark [54, 56], etc.
 Table 4: Overview of Research and Industry Practice in AI Safety Topics
@@
 5.1 Jailbreaking and Red Teaming
-Jailbreaking [57, 163, 222, 265, 273, 277, 295, 314, 323, 335, Zou2023UniversalAT] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs.
+Jailbreaking [58, 167, 228, 272, 280, 284, 302, 324, 333, 345] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs.
 Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content [58, 167, 228, 272, 284, 302, 345].